### v1

# An Evolutionary Algorithm to Accelerate High-Dimensional Optimization

Abstract.
High-dimensional numerical optimization presents a persistent challenge. This paper introduces Quasi-Adaptive Search with Asymptotic Reinitialization (QUASAR), an evolutionary algorithm to accelerate convergence in complex, non-differentiable problems afflicted by the curse of dimensionality.
Evaluated on the notoriously difficult CEC2017 benchmark suite of 29 functions, QUASAR achieved the lowest overall rank sum (150) using the Friedman test, significantly outperforming L-SHADE (229) and standard DE (305) in the dimension-variant trials. QUASAR also proves computationally efficient, with run times averaging $1.4\times$ faster than DE and $7.8\times$ faster than L-SHADE ($p\ll 0.001$) in the population-variant trials.
Building upon Differential Evolution (DE), QUASAR introduces a highly stochastic architecture to dynamically balance exploration and exploitation. Inspired by the probabilistic behavior of quantum particles in a stellar core, the algorithm implements three primary components that augment standard DE mechanisms: 1) probabilistically selected mutation strategies and scaling factors; 2) rank-based crossover rates; 3) asymptotically decaying reinitialization that leverages a covariance matrix of the best solutions to introduce high-quality genetic diversity.
QUASAR’s performance establishes it as an effective, user-friendly optimizer for complex high-dimensional problems.
copyright: none
1. Introduction
High-dimensional numerical optimization remains a persistent challenge across diverse fields, including machine learning, science, engineering, and finance (zhou2024evolutionary). This challenge stems from the curse of dimensionality; the exponential growth of the search space renders gradient-based and parametric search methods ineffective (shan2008survey).
Consequently, optimization metaheuristics, particularly evolutionary algorithms, have become essential for locating optimal solutions in complex function landscapes.
Among these methods, Differential Evolution (DE) has established itself as a robust and highly effective optimizer due to its simplicity, efficiency, and effective mutation strategies (storn1996usage). However, standard DE variants can struggle to maintain an effective balance between global exploration and local exploitation, particularly in high-dimensional search spaces. This often leads to stagnation or premature convergence, requiring frequent restarts or extensive hyperparameter tuning; a significant practical drawback for real-world applications.
To address these limitations, this paper proposes Quasi-Adaptive Search with Asymptotic Reinitialization (QUASAR), an evolutionary algorithm designed to dynamically balance the exploration-exploitation trade-off and achieve superior performance in higher dimensions. This method is conceptually inspired by the search for stability in a stellar core, where quantum particles probabilistically explore configurations searching for the most stable state.
QUASAR extends the DE framework by incorporating a highly stochastic multi-strategy approach, including probabilistically selected mutation strategies and factors, as well as an adaptive crossover rate based on solution fitness. The algorithm’s primary innovation is an asymptotically decaying reinitialization mechanism, utilizing the spatial covariance matrix of top-performing solutions to inject high-quality genetic diversity into the population.
The efficacy of QUASAR is demonstrated through rigorous testing against standard DE and its state-of-the-art variant L-SHADE across the complete CEC2017 benchmark suite. The results show that QUASAR achieves large and statistically significant performance gains, achieving an overall Friedman rank sum of 150, far below L-SHADE (229) or DE (305). These improvements are achieved using default parameters, highlighting QUASAR’s ease of use.
The remainder of this paper is organized as follows: Section 2 details the QUASAR algorithm, including its mutation, crossover, selection, and reinitialization mechanisms. Section 3 defines the experimental design, and Section 4 discusses the experimental results compared to DE and L-SHADE. Section 5 details the ease of use and implementation of the algorithm. Section 6 concludes the work and Section 7 suggests future work.
2. QUASAR Algorithm
Quasi-Adaptive Search with Asymptotic Reinitialization (QUASAR) operates using a highly stochastic, multi-strategy approach.
2.1. Initial Population
The initial population is generated as a Sobol sequence by default, due to its low discrepancy and superior uniformity in high-dimensional spaces compared to other Quasi-Monte Carlo sampling methods (bonneel2025sobol). The default population size, $N$, is set to $10\times D$ (where $D$ is the number of dimensions).
Additional initialization methods are available, including uniform Latin Hypercube and Random distributions, as well as exploitative Hyperellipsoid Density sequences (hyperellipsoid).
2.2. Mutation
Solutions are recombined with one another, analogous to particles sharing information via quantum entanglement. This ’spooky action at a distance’ (A. Einstein, 1947), serves as the conceptual basis for the mutation mechanisms below.
QUASAR uses three distinct mutation strategies, each being a modified Differential Evolution variant. These strategies are selected dynamically based on the specified entanglement probability, entanglerateentangle\ rate. This value defaults to $0.33$, resulting in the three strategies being applied equally across the population.
2.2.1. Local Mutation (Exploitation)
The local mutation strategy, ’Spooky-Best’, is chosen with probability entanglerateentangle\ rate. This strategy is similar to the $\text{DE}/\text{best}/1$ format, exploiting the immediate search space around the current best solution $\mathbf{X}_{\text{best},g}$.
(Spooky-Best) $\mathbf{v}_{i,g}=\mathbf{X}_{\text{best},g}+F_{\text{local}}(\mathbf{X}_{i,g}-\mathbf{X}_{\text{rand},g})$
The associated mutation factor $F_{\text{local}}$ is sampled from a normal distribution $\mathcal{N}(0,0.33^{2})$, allowing for small bi-directional perturbations about the best solution.
2.2.2. Global Mutations (Exploration)
Two global mutation strategies are applied to the remaining (1entanglerate)(1-entangle\ rate) fraction of the population. Promoting global exploration, these strategies, ’Spooky-Current’ and ’Spooky-Random’, each have a 50% probability of being selected:
(Spooky-Current) $\mathbf{v}_{i,g}=\mathbf{X}_{i,g}+F_{\text{global}}(\mathbf{X}_{\text{best},g}-\mathbf{X}_{\text{rand},g})$
(Spooky-Random) $\mathbf{v}_{i,g}=\mathbf{X}_{\text{rand},g}+F_{\text{global}}(\mathbf{X}_{i,g}-\mathbf{X}_{\text{rand},g})$
The mutation factor $F_{\text{global}}$ is sampled from a bimodal distribution $\mathcal{N}(0.5,0.25^{2})+\mathcal{N}(-0.5,0.25^{2})$ to drive non-local exploration. It is worth clarifying that the Spooky-Random strategy uses the same random vector $\mathbf{X}_{\text{rand}}$ twice.
The terms used in the three mutation strategies are defined as:
  • $\mathbf{v}_{i,g}$: The mutant vector for the $i$-th individual at generation $g$.
  • $\mathbf{X}_{i,g}$: The current solution vector.
  • $\mathbf{X}_{\text{best},g}$: The best solution vector known in the population.
  • $\mathbf{X}_{\text{rand},g}$: A randomly chosen solution vector.
  • $F_{\text{local}}$ and $F_{\text{global}}$: The corresponding mutation factors.
Figure 1. Mutation factor distributions for global and local mutation strategies.
2.3. Crossover
QUASAR uses binomial crossover with a dynamic crossover rate ($CR_{i}$), inversely proportional to the solution’s fitness rank. This mechanism ensures that unstable solutions have a higher chance of inheriting components from the new mutant vector $\mathbf{v}_{i}$.
The raw crossover rate ($CR_{raw},i$) for the $i$-th solution is calculated based on its rank ($rank_{i}$), where $rank=1$ corresponds to the best solution (lowest fitness). The final dynamic crossover rate $CR_{i}$ is capped at a minimum of $0.33$:
(1) $CR_{raw,i}=\frac{\text{max\_rank}-\text{rank}_{i}}{\text{max\_rank}}=\frac{N-1-\text{rank}_{i}}{N-1}$
(2) $CR_{i}=\max(CR_{raw,i},0.33)$
The trial vector $\mathbf{u}_{i}$ is then generated component-wise ($n$ dimension) using binomial crossover, comparing a uniform random number $\text{rand}(0,1)$ against $CR_{i}$:
(3) $u_{i,n}=\begin{cases}v_{i,n},&\text{if }\text{rand}(0,1)\leq CR_{i}\\ X_{i,n},&\text{if }\text{rand}(0,1)>CR_{i}\end{cases}$
2.4. Selection
A greedy elitism strategy is used to determine the next generation’s population, where the target vector for generation $g+1$ is the better of the two solutions between the current vector $\mathbf{X}_{i,g}$ and the trial vector $\mathbf{u}_{i,g}$:
(4) $\mathbf{X}_{i,(g+1)}=\begin{cases}\mathbf{u}_{i},&\text{if }f(\mathbf{u}_{i})<f(\mathbf{X}_{i})\\ \mathbf{X}_{i},&\text{if }f(\mathbf{u}_{i})\geq f(\mathbf{X}_{i})\end{cases}$
2.5. Asymptotic Reinitialization
The covariance-guided vector reinitialization applies to a fixed 33% of the worst-performing solutions. These solutions have a high chance in early generations of ’tunneling’ to more promising regions.
The probability of reinitialization $P_{\text{reinit}}(g)$ decays asymptotically from $P_{\text{reinit}}(0)=1$ toward a target final probability, calculated using the current generation $g$:
(5) $P_{\text{reinit}}(g)=\exp\left(\frac{\ln(P_{{}_{\text{final}}})}{g_{{}_{\text{final}}}g_{{}_{\text{max}}}}\cdot g\right)$
where $g_{\max}$ is the maximum generation count, $P_{\text{final}}=0.33$ is the target final probability, and $g_{final}=0.33$ defines the fraction of generations at which $P_{reinit}$ reaches $P_{final}$. This ensures the probability drops to $P_{\text{final}}$ at $g=g_{\max}\cdot g_{\text{final}}$.
When a reinitialization occurs, new positions are sampled, informed by the covariance of top-performing individuals. The best $M$ solutions (where $M=25\%\text{ of the population}$) are used to calculate the mean vector $\boldsymbol{\mu}$ and the covariance matrix $\boldsymbol{\Sigma}$:
(6) $\boldsymbol{\mu}=\frac{1}{M}\sum_{i=1}^{M}\mathbf{x}_{i}$
(7) $\boldsymbol{\Sigma}=\frac{1}{M-1}\sum_{i=1}^{M}(\mathbf{x}_{i}-\boldsymbol{\mu})(\mathbf{x}_{i}-\boldsymbol{\mu})^{\text{T}}+\epsilon\mathbf{I}$
where $\epsilon$ is a small scalar ($10^{-12}$) and $\mathbf{I}$ is the identity matrix, added to prevent $\boldsymbol{\Sigma}$ from being singular.
New vectors $\mathbf{y}$ are generated by sampling from a multivariate normal distribution, with noise added to further promote exploration.
(8) $\mathbf{y}\sim\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})+\boldsymbol{\delta}$
The noise vector is $\boldsymbol{\delta}$ is sampled from a multivariate normal distribution, with respect to the search bounds:
(9) $\boldsymbol{\delta}\sim\mathcal{N}\left(\mathbf{0},\left(\frac{\mathbf{b_{high,n}}-\mathbf{b_{low,n}}}{20}\right)^{2}\mathbf{I}\right)$
Finally, all components of the new solution $\mathbf{X}_{\text{new}}$ are clipped to ensure they remain within the defined search bounds $\mathbf{B}=[\mathbf{b}_{\text{low}},\mathbf{b}_{\text{high}}]$.
(10) $\mathbf{X_{new,n}}=\text{clip}(\mathbf{y_{n}},\mathbf{B_{n}})$
These new solutions completely replace the reinitialized solutions and are not subjected to crossover mechanisms.
Due to the $\mathcal{O}(N)$ scaling of the covariance matrix calculations, along with decaying probability of occurrence, the additional overhead becomes negligible compared to the objective function evaluations.
Figure 2. QUASAR workflow for evolving each generation.
3. Experimental Design
3.1. Comparison Optimizers & Test Functions
The performance of QUASAR was benchmarked against modern implementations of Differential Evolution (DE) and L-SHADE, specifically SciPy’s DE and MealPy’s L-SHADE.
This choice is justified by the fact that the QUASAR algorithm is built on the DE framework, as well as DE’s standing as a robust and effective optimizer (price1996differential). L-SHADE was specifically chosen due to its success in CEC optimization competitions (mo2019efficient).
The trials were conducted using all 29 functions of the CEC2017 numerical optimization benchmark suite.
3.2. Experimental Setup
To ensure a fair comparison, each optimizer was run using their default hyperparameters. The initial population generation methods were also left to defaults; the tested sample sizes not being powers of 2 diminishes the uniformity of QUASAR’s initial Sobol sequence (joe2008constructing), providing a theoretical initial handicap.
Since SciPy’s DE implementation internally scales the input population size by dimensionality ($popsize\times D$ when all dimensions have equal bounds), a correction factor of $1/D$ is applied to its initial sample size.
The computational budget for metaheuristic comparison is often defined by the maximum number of objective function evaluations ($FES_{max}$) (kazikova2021objective). Due to L-SHADE’s linear population reduction, the $FES_{max}$ becomes a difficult comparison metric to compare against. As such, the primary comparisons are performed on final solution fitness errors and overall optimization times, given the same number of generations, using modern implementations via SciPy’s DE and MealPy’s L-SHADE.
By default, the L-SHADE implementation features tolerance parameters for early convergence, and the QUASAR and DE implementations further include a final polishing step using L-BFGS-B minimization. These were all disabled for the trials to ensure the optimizers fully utilize their computational budget, terminating only when reaching the maximum number of generations $g_{max}$.
3.3. Experimental Trials:
The trials were structured into two distinct phases to assess QUASAR’s behavior under varying dimensionalities and population sizes.
The results are compared based on the overall aggregated rank sums, geometric mean error factors ($GMERF$), and optimization run times.
Dimension-Variant Trials:
These trials assessed QUASAR’s scalability by varying the problem dimension $D$ with a fixed population size ($N=1000$).
The dimension-variant experiments were run with $30$ independent trials for each function/dimension combination to ensure statistical significance.
The dimensions were varied between $[10,30,50,100]$, resulting in a total of 10,400 full optimization runs; 3,480 for each algorithm. The computational budget was kept constant, with maximum number of iterations $g_{max}=100$ and sample size $N=1000$.
Table 1. Dimension-Variant Test Parameters
Parameter Test Values
Dimensions ($D$) [10, 30, 50, 100]
Population Size ($N$) 1000
Generations ($g_{\max}$) 100
Sample-Variant Trials:
These trials assessed the impact of population size on performance in high dimensions (a fixed $D=100$). The computational budget is varied between population sizes of [100, 250, 500, 1000, 5000], with 30 independent trials performed for each sample size / test function combination using maximum number of iterations $g_{max}=100$.
The $N=5000$ trials, however, were only run for 10 trials to ensure statistical significance while mitigating the high computational overhead, resulting in a total of 11,310 full optimization runs; 3770 for each algorithm.
Table 2. Sample-Variant Test Parameters
Parameter Test Values
Dimensions ($D$) 100
Population Size ($N$) [100, 250, 500, 1000, 5000]
Generations ($g_{\max}$) 100
3.4. Statistical Analysis:
Overall Rank:
To draw the overall statistical comparison between the optimizers across all function/dimension and function/sample size combinations, the Friedman test was used.
For each unique combination, the median final error value ($f$) across all independent trials was calculated for each optimizer. Optimizers were ranked based on these median errors, with the best performance (lowest error) receiving a rank of 1.
The ranks were summed across all scenarios to produce the aggregated rank sum for each optimizer. A lower sum indicates superior overall performance.
Solution Quality Improvements:
The primary comparison metric used to quantify the magnitude of performance difference is the Geometric Mean Error Reduction Factor $GMERF$, which provides an unbiased measure of relative performance across all $N_{\text{trials}}$ independent runs for a given scenario:
$\text{GMERF}=\left(\prod_{i=1}^{N_{\text{trials}}}\frac{f_{\text{comparison},i}}{f_{\text{QUASAR},i}}\right)^{\frac{1}{N_{\text{trials}}}}$
Here, $f_{\text{comparison},i}$ and $f_{\text{QUASAR},i}$ are the final error values achieved by the competitor algorithm (e.g., DE or L-SHADE) and QUASAR, respectively, in the $i$-th trial. GMERF>1{GMERF}>1 indicates that QUASAR performed better than the comparison optimizer.
For the final aggregated results, the Overall $GMERF$ is calculated as the geometric mean of the GMERF values obtained for each function/dimension combination.
$\text{GMERF}_{\text{Overall}}=\left(\prod_{j=1}^{N_{\text{combinations}}}\text{GMERF}_{j}\right)^{\frac{1}{N_{\text{combinations}}}}$
Computational Efficiency:
The computational efficiency of the optimizers is quantified by comparing their average optimization run times.
The ratios of run times $R_{t}$ for QUASAR vs. DE & L-SHADE are shown in Table 7 & Table 8, by dimension and sample size respectively. $R_{t}$ is calculated by dividing the average run time of the comparison method ($\text{Mean Sec}_{\text{comparison}}$) by the average run time of QUASAR ($\text{Mean Sec}_{\text{QUASAR}}$) across all trials for that specific dimension.
(11) $R_{t}=\frac{\text{Mean Run Time}_{\text{comparison},D}}{\text{Mean Run Time}_{\text{QUASAR},D}}$
The overall efficiency metrics are the overall run time ratios, $R_{t_{overall}}$, calculated as the arithmetic mean of each paired ratio $R_{t}$.
(12) $\text{R}_{\text{t,overall}}=\frac{1}{|D|}\sum_{D\in D}\left(\frac{1}{|R_{t,D}|}\sum_{i=1}^{|R_{t,D}|}\frac{t_{\text{comparison},i,D}}{t_{\text{QUASAR},i,D}}\right)$
Statistical Significance:
For direct comparison within each dimension or sample size, the paired Wilcoxon signed-rank test is used. This non-parametric test was chosen as the final error distributions are on different scales for each function and are assumed not to be normal, and to ensure consistency with the non-parametric ranking methodology.
The Wilcoxon test was conducted on the final error values ($f$) and run times to generate the p-values for each function/dimension combination.
4. Results
The primary solution quality metrics are the Friedman aggregated rank sum and the geometric mean error reduction factor $GMERF$, quantifying QUASAR’s improvements in final solution fitness. The computational efficiency is also measured by comparing full optimization run times.
Since lower errors and run times indicate better performance, a $\text{ratio}>1$, for both the $GMERF$ and run times ratios, indicates that QUASAR demonstrated superior performance.
Similarly, all plots shown are of raw fitness errors and run times; for all plots, lower values indicate higher performance.
4.1. Solution Quality
The ranks and geometric means of fitness error $GMERF$ of the final solution are calculated for both the dimension-variant and sample-variant trials.
The 95% confidence intervals show consistent performance improvements in almost every scenario tested.
4.1.1. Dimension-Variant Trials:
The dimension-variant analysis assessed QUASAR’s convergence against DE and L-SHADE across four discrete dimensions ($10D$, $30D$, $50D$, and $100D$), using a fixed sample size $N=\text{1000}$ and maximum generations $g_{max}=100$.
The geometric mean error reduction factors ($GMERF$) were consistently high across all dimensions (Fig 3). The 95% confidence intervals validate its consistent improvements; the intervals do not fall below $1.0\times$ GMERFGMERF.
The largest improvements were seen against DE in $100D$, achieving a $GMERF$ of $\text{13.52}\times$ with a CI upper bound of $16.68\times$. The comparison against L-SHADE showed similar improvements; the full set of $GMERF$ values are shown in Table 5.
Figure 3. Geometric mean error (log-scale) by dimension for each optimizer, with population size of 1000.
The Friedman test was used on the entire set of unique function/dimension combinations to determine the overall performance hierarchy of the optimizers. The test confirmed a statistically significant improvement in performance ($p\ll 0.001$), shown in Table 3.
Table 3. Dimension-Variant Rank Sums
Method Rank Sum
QUASAR 150
L-SHADE 229
DE 305
4.1.2. Sample Size-Variant Trials:
To assess the impact of varying population sizes, QUASAR and DE were compared across a fixed $100D$ and 100 maximum generations $g_{max}$, using various sample sizes ($N$).
The fitness improvement factors saw the highest gains with higher population sizes; in the $N=5000$ case, QUASAR reached a geometric mean error improvement factor $GMERF$ of $\text{18.5}\times$ compared to DE, and $4.47\times$ compared to L-SHADE.
These performance increases scale super-linearly with population size (Fig. 4, Table 6). The only scenario where QUASAR did not outperform L-SHADE was the $N=100$ (low-population) case, where L-SHADE achieved a $GMERF$ of $0.49\times$.
The Friedman test is run on the set of all function/population size combinations, shown in Table 4.
Figure 4. Geometric mean error (log-scale) by population size for each optimizer, in 100 dimensions.
Table 4. Sample-Variant Rank Sums
Method Rank Sum
QUASAR 217
L-SHADE 223
DE 430
Table 5. Dimension-Variant GMERF Results: QUASAR vs L-SHADE and DE
D vs L-SHADE vs DE p-value p-value QUASAR L-SHADE DE
GMERF 95% CI GMERF 95% CI (vs L-SHADE) (vs DE) (GM Error) (GM Error) (GM Error)
10 1.30 (1.20, 1.40) 1.06 (1.01, 1.12) $7\times 10^{-16}$ $7\times 10^{-3}$ $1.59\times 10^{3}$ $2.10\times 10^{3}$ $1.72\times 10^{3}$
30 4.14 (3.62, 4.74) 4.94 (4.24, 5.77) $5\times 10^{-91}$ $3\times 10^{-108}$ $5.10\times 10^{3}$ $2.11\times 10^{4}$ $2.52\times 10^{4}$
50 3.78 (3.32, 4.31) 8.44 (7.09, 10.04) $9\times 10^{-79}$ $5\times 10^{-134}$ $1.46\times 10^{4}$ $5.51\times 10^{4}$ $1.23\times 10^{5}$
100 2.67 (2.32, 3.06) 13.52 (10.95, 16.68) $2\times 10^{-36}$ $2\times 10^{-141}$ $7.917\times 10^{4}$ $2.11\times 10^{5}$ $1.07\times 10^{6}$
Table 6. Sample Size-Variant GMERF Results: QUASAR vs L-SHADE & DE
S vs L-SHADE vs DE p-value p-value QUASAR L-SHADE DE
GMERF 95% CI GMERF 95% CI (vs L-SHADE) (vs DE) (GM Error) (GM Error) (GM Error)
100 0.49 (0.44, 0.56) 2.43 (2.04, 2.89) $7\times 10^{-86}$ $3\times 10^{-46}$ $5.94\times 10^{5}$ $2.94\times 10^{5}$ $1.44\times 10^{6}$
250 1.34 (1.19, 1.50) 7.86 (6.51, 9.49) $3.3\times 10^{-2}$ $3\times 10^{-141}$ $1.79\times 10^{5}$ $2.39\times 10^{5}$ $1.41\times 10^{6}$
500 1.93 (1.70, 2.20) 10.40 (8.54, 12.67) $2\times 10^{-8}$ $2\times 10^{-140}$ $1.15\times 10^{5}$ $2.21\times 10^{5}$ $1.19\times 10^{6}$
1000 2.66 (2.32, 3.06) 13.52 (10.95, 16.68) $5\times 10^{-34}$ $2\times 10^{-141}$ $7.92\times 10^{4}$ $2.11\times 10^{5}$ $1.07\times 10^{6}$
5000 4.47 (3.41, 5.86) 18.54 (12.49, 27.53) $3\times 10^{-37}$ $3\times 10^{-48}$ $4.20\times 10^{4}$ $1.88\times 10^{5}$ $7.78\times 10^{5}$
4.2. Computational Efficiency
The computational efficiency of QUASAR was measured using the overall ratio of run time $R_{t_{overall}}$ (seconds) relative to DE and L-SHADE.
4.2.1. Dimension-Variant Trials:
In the dimension-variant trials, QUASAR demonstrated consistently faster run times compared to DE and L-SHADE.
The overall average run time improvement ratios $R_{t_{overall}}$ were $1.42\times$ and $6.29\times$ respectively. The average paired run time ratios $R_{t}$ are shown in Table 7, and optimization run times in seconds are in Table 7 and Fig. 5.
Figure 5. Average optimization run times for each optimizer, by dimension.
4.2.2. Sample Size-Variant Trials:
In the sample-variant trials, run times were consistently faster than DE and L-SHADE, averaging $7.78\times$ and $1.39\times$ speed increases respectively across all sample-variant trials (Fig. 6, Table 8).
The $N=250$ trials, however, demonstrated a ratio $R_{t}$ vs DE of $0.84\times$, indicating that DE was approximately 16% faster.
Figure 6. Average optimization run times for each optimizer, by population size.
Table 7. Dimension-Variant Run Times: QUASAR vs L-SHADE & DE
D $\mathbf{R_{t}}$ $\mathbf{R_{t}}$ p-value p-value QUASAR L-SHADE DE
$\mathbf{vs\ \textbf{L-SHADE}}$ $\mathbf{vs\ DE}$ (vs L-SHADE) (vs DE) (sec) (sec) (sec)
10 $4.21$ $1.26$ $5\text{\times}{10}^{-144}$ $7.3\text{\times}{10}^{-141}$ $16.4$ $69.1$ $20.7$
30 $3.17$ $1.17$ $5\text{\times}{10}^{-144}$ $1.1\text{\times}{10}^{-139}$ $25.0$ $79.4$ $29.2$
50 $2.66$ $1.14$ $5\text{\times}{10}^{-144}$ $1.6\text{\times}{10}^{-137}$ $33.5$ $89.0$ $38.1$
100 $2.03$ $1.09$ $5\text{\times}{10}^{-144}$ $3\text{\times}{10}^{-137}$ $55.4$ $112.4$ $60.5$
Table 8. Sample Size-Variant Run Times: QUASAR vs L-SHADE & DE
S $\mathbf{R_{t}}$ $\mathbf{R_{t}}$ p-value p-value QUASAR L-SHADE DE
$\mathbf{vs\ \textbf{L-SHADE}}$ $\mathbf{vs\ DE}$ (vs L-SHADE) (vs DE) (sec) (sec) (sec)
100 $1.37$ $1.05$ $\text{\times}{10}^{-142} 1×1$$\text{\times}{10}^{-84}$ $5.8$ $7.9$ $6.0$
250 $1.47$ $0.84$ $1\text{\times}{10}^{-142}$ $5\text{\times}{10}^{-58}$ $13.7$ $20.1$ $11.4$
500 $1.65$ $1.06$ $1\text{\times}{10}^{-142}$ $2\text{\times}{10}^{-123}$ $28.7$ $47.2$ $30.5$
1000 $2.00$ $1.08$ $1\text{\times}{10}^{-142}$ $9\text{\times}{10}^{-121}$ $57.3$ $114.4$ $62.0$
5000 $5.96$ $1.29$ $3\text{\times}{10}^{-49}$ $3\text{\times}{10}^{-48}$ $292.0$ $1741.7$ $376.1$
5. Ease of Use and Implementation
QUASAR is designed to be robust with minimal parameter dependence, making it effective for most numerical optimization problems.
5.1. Ease of Use
A key feature of QUASAR is its ability to operate effectively with its default entanglerateentangle\ rate, bypassing the hyperparameter tuning often required by other metaheuristics such as DE and L-SHADE.
The algorithm relies on its quasi-adaptive mechanisms to dynamically balance exploration and exploitation. For typical use cases, the only inputs needed are the objective function, search bounds, and the target number of generations ($g_{\max}$).
5.2. Implementation
QUASAR is implemented as a streamlined open-source Python package, designed for seamless integration into existing workflows.
The source code requires minimal dependencies, relying only on NumPy and SciPy. The open-source package, hdim-opt, is available at github.com/jgsoltes/hdim-opt for full use and reproducibility.
The package follows the syntax structures of SciPy’s optimization module, making it nearly one-to-one interchangeable with its Differential Evolution function.
6. Conclusion
This study introduced Quasi-Adaptive Search with Asymptotic Reinitialization (QUASAR), a quantum-inspired evolutionary algorithm designed to efficiently explore and exploit high-dimensional spaces.
The algorithm utilizes a highly stochastic architecture, including probabilistic mutation mechanisms, rank-based crossover rates, and asymptotically decaying covariance reinitializations.
Experimental trials across the CEC2017 benchmark suite confirm that QUASAR delivers exceptional performance compared to Differential Evolution (DE) and L-SHADE, consistently ranking best via the Friedman test.
[…]
The computational efficiency of the algorithm is a key benefit, with run times consistently lower than the comparison optimizers ($R_{t}>1$), scaling positively with population size and dimensionality.
The empirical evidence validates QUASAR as a powerful evolutionary algorithm, ready for the next generation of high-dimensional optimization problems.
7. Future Works
7.1. Enhancements:
QUASAR’s quasi-adaptive mechanisms are highly effective; future improvements will explore fully self-adaptive strategies, similar to those found in other modern DE variants.
The current covariance matrix ($\mathbf{\Sigma}$) treats all top solutions equall;. future work will investigate applying weights to the very best solutions. New methods for defining the probability of occurrence may be explored, such as sinusoidal or sigmoid activation behavior.
Crossover rate is currently defined by fitness rank, and may benefit from a probability-based approach similar to the other QUASAR components. Soft-max probabilities or similar methods to determine recombination probabilities may enhance genetic diversity.
7.2. Verification:
QUASAR’s highly stochastic framework makes it a practical choice for non-differentiable and non-parametric problems. The algorithms’s robustness would be further validated by rigorously evaluating its performance on more constrained, non-parametric problems.
Acknowledgements.
The author would like to thank Dr. Kellen Sorauf, Dr. Ksenia Polson, Dr. Mike Busch, and Monet Morris for their invaluable feedback and support throughout this study. It would not have been possible without the insights gained from each of our discussions. Dr. Sorauf’s recommendation to include L-SHADE in the experiment significantly improved the rigor of the study; Dr. Polson’s lectures laid the framework for the statistical analysis; Dr. Busch pushed me to pursue the project to begin with. Monet remains a source of inspiration, as always. This research was conducted independently and was not part of a formal project.

### v2

@@
 5.2. Implementation
-QUASAR is implemented as a streamlined open-source Python package, designed for seamless integration into existing workflows.
-The source code requires minimal dependencies, relying only on NumPy and SciPy. The open-source package, hdim-opt, is available at github.com/jgsoltes/hdim-opt for full use and reproducibility.
-The package follows the syntax structures of SciPy’s optimization module, making it nearly one-to-one interchangeable with its Differential Evolution function.
+QUASAR is implemented as part of a streamlined open-source Python package, designed for seamless integration into existing workflows.
+The syntax follows the structure of SciPy’s optimization module, making it nearly one-to-one interchangeable with its Differential Evolution.
+The source code requires minimal dependencies, relying only on NumPy and SciPy.
+The open-source package containing QUASAR, hdim-opt, is available via the Python Package Index (PyPI) using ’pip install hdim_opt’, or at github.com/jgsoltes/hdim-opt which includes the experimental notebooks for full reproducibility.
 6. Conclusion
@@
 New methods for defining the probability of occurrence may be explored, such as sinusoidal or sigmoid activation behavior.
-Crossover rate is currently defined by fitness rank, and may benefit from a probability-based approach similar to the other QUASAR components.
-Soft-max probabilities or similar methods to determine recombination probabilities may enhance genetic diversity.
+Crossover rate is currently defined by fitness rank, and may benefit from a probability-based approach, similar to the other QUASAR components, via soft-max or similar methods.
 7.2. Verification:

### v3

@@
 Abstract.
-High-dimensional numerical optimization presents a persistent challenge.
+High-dimensional numerical optimization presents a persistent challenge in computational science.
 This paper introduces Quasi-Adaptive Search with Asymptotic Reinitialization (QUASAR), an evolutionary algorithm to accelerate convergence in complex, non-differentiable problems afflicted by the curse of dimensionality.
-Evaluated on the notoriously difficult CEC2017 benchmark suite of 29 functions, QUASAR achieved the lowest overall rank sum (150) using the Friedman test, significantly outperforming L-SHADE (229) and standard DE (305) in the dimension-variant trials.
-QUASAR also proves computationally efficient, with run times averaging $1.4\times$ faster than DE and $7.8\times$ faster than L-SHADE ($p\ll 0.001$) in the population-variant trials.
-Building upon Differential Evolution (DE), QUASAR introduces a highly stochastic architecture to dynamically balance exploration and exploitation.
-Inspired by the probabilistic behavior of quantum particles in a stellar core, the algorithm implements three primary components that augment standard DE mechanisms: 1) probabilistically selected mutation strategies and scaling factors; 2) rank-based crossover rates; 3) asymptotically decaying reinitialization that leverages a covariance matrix of the best solutions to introduce high-quality genetic diversity.
-QUASAR’s performance establishes it as an effective, user-friendly optimizer for complex high-dimensional problems.
+QUASAR expands upon the core principles of Differential Evolution (DE), introducing quasi-adaptive mechanisms to dynamically balance exploration and exploitation in its search.
+Inspired by the behavior of quantum particles, the algorithm utilizes three highly stochastic mechanisms that augment standard DE: 1) probabilistic mutation strategies and scaling factors; 2) rank-based crossover rates; 3) asymptotically decaying covariance reinitializations.
+Evaluated on the notoriously difficult CEC2017 benchmark suite of 29 test functions, QUASAR achieved the lowest overall rank sum (367) using the Friedman test, outperforming DE (735) and L-SHADE (452).
+Geometric mean comparisons show average final solution quality improvements of $3.85\times$ and $2.07\times$ compared to DE and L-SHADE, respectively ($p\ll 0.001$), with average optimization speed averaging $1.40\times$ and $5.16\times$ faster.
+QUASAR’s performance establishes it as an effective, efficient, and user-friendly evolutionary algorithm for complex high-dimensional problems.
 copyright: none
 1. Introduction
-High-dimensional numerical optimization remains a persistent challenge across diverse fields, including machine learning, science, engineering, and finance (zhou2024evolutionary).
-This challenge stems from the curse of dimensionality; the exponential growth of the search space renders gradient-based and parametric search methods ineffective (shan2008survey).
+High-dimensional numerical optimization remains a persistent challenge across many fields, including machine learning, science, engineering, and finance (Zhou et al., 2024).
+This challenge stems from the curse of dimensionality; the exponential growth of the search space renders gradient-based and parametric search methods ineffective (Shan and Wang, 2008).
 Consequently, optimization metaheuristics, particularly evolutionary algorithms, have become essential for locating optimal solutions in complex function landscapes.
-Among these methods, Differential Evolution (DE) has established itself as a robust and highly effective optimizer due to its simplicity, efficiency, and effective mutation strategies (storn1996usage).
+Among these methods, Differential Evolution (DE) has established itself as a robust and highly effective optimizer due to its simplicity, efficiency, and effective mutation strategies (Storn, 1996).
 However, standard DE variants can struggle to maintain an effective balance between global exploration and local exploitation, particularly in high-dimensional search spaces.
-This often leads to stagnation or premature convergence, requiring frequent restarts or extensive hyperparameter tuning; a significant practical drawback for real-world applications.
-To address these limitations, this paper proposes Quasi-Adaptive Search with Asymptotic Reinitialization (QUASAR), an evolutionary algorithm designed to dynamically balance the exploration-exploitation trade-off and achieve superior performance in higher dimensions.
+This often leads to stagnation or premature convergence that requires frequent restarts or extensive hyperparameter tuning, a significant practical drawback for real-world applications.
+To address these limitations, this paper proposes Quasi-Adaptive Search with Asymptotic Reinitialization (QUASAR), an evolutionary algorithm designed to dynamically balance the exploration-exploitation trade-off and achieve superior performance in high-dimensional spaces.
 This method is conceptually inspired by the search for stability in a stellar core, where quantum particles probabilistically explore configurations searching for the most stable state.
@@
 The efficacy of QUASAR is demonstrated through rigorous testing against standard DE and its state-of-the-art variant L-SHADE across the complete CEC2017 benchmark suite.
-The results show that QUASAR achieves large and statistically significant performance gains, achieving an overall Friedman rank sum of 150, far below L-SHADE (229) or DE (305).
-These improvements are achieved using default parameters, highlighting QUASAR’s ease of use.
+The results show large and statistically significant performance gains; these improvements are achieved using default parameters, highlighting QUASAR’s ease of use.
 The remainder of this paper is organized as follows: Section 2 details the QUASAR algorithm, including its mutation, crossover, selection, and reinitialization mechanisms.
-Section 3 defines the experimental design, and Section 4 discusses the experimental results compared to DE and L-SHADE.
-Section 5 details the ease of use and implementation of the algorithm.
-Section 6 concludes the work and Section 7 suggests future work.
+Section 3 defines the experimental design, and Section 4 analyzes the experimental results compared to DE and L-SHADE.
+Section 5 describes the ease of use and implementation of the algorithm.
+Section 6 concludes the paper, with Section 7 proposing future work.
 2. QUASAR Algorithm
 Quasi-Adaptive Search with Asymptotic Reinitialization (QUASAR) operates using a highly stochastic, multi-strategy approach.
+The mechanisms defined below synergize to create a dynamic ’push-and-pull’ effect on the solution vectors.
 2.1. Initial Population
-The initial population is generated as a Sobol sequence by default, due to its low discrepancy and superior uniformity in high-dimensional spaces compared to other Quasi-Monte Carlo sampling methods (bonneel2025sobol).
-The default population size, $N$, is set to $10\times D$ (where $D$ is the number of dimensions).
-Additional initialization methods are available, including uniform Latin Hypercube and Random distributions, as well as exploitative Hyperellipsoid Density sequences (hyperellipsoid).
+The initial population is generated as a Sobol sequence by default, due to its low discrepancy and superior uniformity in high-dimensional spaces compared to other quasi-Monte Carlo sampling methods (Bonneel et al., 2025).
+The default population size, $N$, is set to $10\times D$, where $D$ is the number of dimensions.
+Additional initialization methods include uniform Latin Hypercube and random distributions, as well as non-uniform Hyperellipsoid Density sequences (Soltes, 2025).
 2.2. Mutation
@@
 This ’spooky action at a distance’ (A.
-Einstein, 1947), serves as the conceptual basis for the mutation mechanisms below.
+Einstein, 1947) serves as the conceptual basis for the mutation mechanisms below.
 QUASAR uses three distinct mutation strategies, each being a modified Differential Evolution variant.
-These strategies are selected dynamically based on the specified entanglement probability, entanglerateentangle\ rate.
+These strategies are selected dynamically based on the specified entanglement probability, entangle rate.
 This value defaults to $0.33$, resulting in the three strategies being applied equally across the population.
-2.2.1. Local Mutation (Exploitation)
-The local mutation strategy, ’Spooky-Best’, is chosen with probability entanglerateentangle\ rate.
-This strategy is similar to the $\text{DE}/\text{best}/1$ format, exploiting the immediate search space around the current best solution $\mathbf{X}_{\text{best},g}$.
+2.2.1. Local Mutation (Exploitation):
+The local mutation strategy, ’Spooky-Best’, is chosen with probability entangle rate.
+This strategy is similar to the DE/best/1 format, exploiting the immediate search space around the current best solution $\mathbf{X}_{\text{best},g}$.
 (Spooky-Best) $\mathbf{v}_{i,g}=\mathbf{X}_{\text{best},g}+F_{\text{local}}(\mathbf{X}_{i,g}-\mathbf{X}_{\text{rand},g})$
 The associated mutation factor $F_{\text{local}}$ is sampled from a normal distribution $\mathcal{N}(0,0.33^{2})$, allowing for small bi-directional perturbations about the best solution.
-2.2.2. Global Mutations (Exploration)
-Two global mutation strategies are applied to the remaining (1entanglerate)(1-entangle\ rate) fraction of the population.
+2.2.2. Global Mutations (Exploration):
+Two global mutation strategies are applied to the remaining $(1-\textit{entangle rate})$ fraction of the population.
 Promoting global exploration, these strategies, ’Spooky-Current’ and ’Spooky-Random’, each have a 50% probability of being selected:
@@
 The terms used in the three mutation strategies are defined as:
-  • $\mathbf{v}_{i,g}$: The mutant vector for the $i$-th individual at generation $g$.
-  • $\mathbf{X}_{i,g}$: The current solution vector.
-  • $\mathbf{X}_{\text{best},g}$: The best solution vector known in the population.
-  • $\mathbf{X}_{\text{rand},g}$: A randomly chosen solution vector.
-  • $F_{\text{local}}$ and $F_{\text{global}}$: The corresponding mutation factors.
+  • $\mathbf{v}_{i,g}$: Mutant vector for the $i$-th individual at generation $g$.
+  • $\mathbf{X}_{i,g}$: Current solution vector.
+  • $\mathbf{X}_{\text{best},g}$: Best known solution in the population.
+  • $\mathbf{X}_{\text{rand},g}$: Randomly chosen solution vector.
+  • $F_{\text{local}}$ and $F_{\text{global}}$: Corresponding mutation factors.
 Figure 1. Mutation factor distributions for global and local mutation strategies.
 2.3. Crossover
-QUASAR uses binomial crossover with a dynamic crossover rate ($CR_{i}$), inversely proportional to the solution’s fitness rank.
-This mechanism ensures that unstable solutions have a higher chance of inheriting components from the new mutant vector $\mathbf{v}_{i}$.
-The raw crossover rate ($CR_{raw},i$) for the $i$-th solution is calculated based on its rank ($rank_{i}$), where $rank=1$ corresponds to the best solution (lowest fitness).
-The final dynamic crossover rate $CR_{i}$ is capped at a minimum of $0.33$:
-(1) $CR_{raw,i}=\frac{\text{max\_rank}-\text{rank}_{i}}{\text{max\_rank}}=\frac{N-1-\text{rank}_{i}}{N-1}$
-(2) $CR_{i}=\max(CR_{raw,i},0.33)$
+QUASAR uses binomial crossover with a dynamic crossover rate ($CR_{i}$) proportional to the solution’s fitness rank.
+The best solutions are rapidly recombined to maximize both exploration and exploitation via synergy with global and local mutations.
+The raw crossover rate ($CR_{\text{raw}},i$) for the $i$-th solution is calculated based on its rank ($\textit{rank}_{i}\in[0,N-1]$), where $\textit{rank}=0$ corresponds to the best solution (lowest objective function value).
+The final dynamic crossover rate $CR_{i}$ is capped at a minimum of $0.33$.
+This mechanism ensures that the best solutions have a higher chance of inheriting components from the new mutant vector $\mathbf{v}_{i}$.
+(1) $CR_{\text{raw},i}=\frac{N-1-\text{rank}_{i}}{N-1}$
+(2) $CR_{i}=\max(CR_{\text{raw},i},0.33)$
 The trial vector $\mathbf{u}_{i}$ is then generated component-wise ($n$ dimension) using binomial crossover, comparing a uniform random number $\text{rand}(0,1)$ against $CR_{i}$:
@@
 2.4. Selection
-A greedy elitism strategy is used to determine the next generation’s population, where the target vector for generation $g+1$ is the better of the two solutions between the current vector $\mathbf{X}_{i,g}$ and the trial vector $\mathbf{u}_{i,g}$:
+A greedy elitism strategy is used to determine the next generation’s population, where the target vector for generation $g+1$ is the better of the two solutions between the current vector $\mathbf{X}_{i,g}$ and the trial vector $\mathbf{u}_{i,g}$.
+If a solution undergoes reinitialization, it bypasses the selection step.
 (4) $\mathbf{X}_{i,(g+1)}=\begin{cases}\mathbf{u}_{i},&\text{if }f(\mathbf{u}_{i})<f(\mathbf{X}_{i})\\ \mathbf{X}_{i},&\text{if }f(\mathbf{u}_{i})\geq f(\mathbf{X}_{i})\end{cases}$
@@
 The covariance-guided vector reinitialization applies to a fixed 33% of the worst-performing solutions.
-These solutions have a high chance in early generations of ’tunneling’ to more promising regions.
+In early generations, these solutions have a high probability of ’tunneling’ to more promising regions.
+This mechanism complements the rank-based crossover; while the best solutions recombine frequently, lower quality solutions explore locally until stochastically tunneling to a better location.
 The probability of reinitialization $P_{\text{reinit}}(g)$ decays asymptotically from $P_{\text{reinit}}(0)=1$ toward a target final probability, calculated using the current generation $g$:
-(5) $P_{\text{reinit}}(g)=\exp\left(\frac{\ln(P_{{}_{\text{final}}})}{g_{{}_{\text{final}}}g_{{}_{\text{max}}}}\cdot g\right)$
-where $g_{\max}$ is the maximum generation count, $P_{\text{final}}=0.33$ is the target final probability, and $g_{final}=0.33$ defines the fraction of generations at which $P_{reinit}$ reaches $P_{final}$.
-This ensures the probability drops to $P_{\text{final}}$ at $g=g_{\max}\cdot g_{\text{final}}$.
-When a reinitialization occurs, new positions are sampled, informed by the covariance of top-performing individuals.
-The best $M$ solutions (where $M=25\%\text{ of the population}$) are used to calculate the mean vector $\boldsymbol{\mu}$ and the covariance matrix $\boldsymbol{\Sigma}$:
+(5) $P_{\text{reinit}}(g)=\exp\left(\frac{\ln(P_{\text{final}})}{g_{\text{final}}\cdot g_{\text{max}}}\cdot g\right)$
+where $g_{\text{max}}$ is the maximum generation count, $P_{\text{final}}=0.33$ is the target final probability, and $g_{\text{final}}=0.33$ defines the fraction of generations at which $P_{reinit}$ reaches $P_{final}$.
+This ensures the probability drops to $P_{\text{final}}$ at $g=g_{\text{max}}\cdot g_{\text{final}}$.
+During a reinitialization event, new solutions are sampled from a multivariate distribution shaped by top performing individuals.
+The top $M$ solutions (where $M=25\%\text{of the population}$) are used to calculate the mean vector $\boldsymbol{\mu}$ and the covariance matrix $\boldsymbol{\Sigma}$:
 (6) $\boldsymbol{\mu}=\frac{1}{M}\sum_{i=1}^{M}\mathbf{X}_{i}$
-(7) $\boldsymbol{\Sigma}=\frac{1}{M-1}\sum_{i=1}^{M}(\mathbf{x}_{i}-\boldsymbol{\mu})(\mathbf{x}_{i}-\boldsymbol{\mu})^{\text{T}}+\epsilon\mathbf{I}$
+(7) $\boldsymbol{\Sigma}=\frac{1}{M-1}\sum_{i=1}^{M}(\mathbf{X}_{i}-\boldsymbol{\mu})(\mathbf{X}_{i}-\boldsymbol{\mu})^{\top}+\epsilon\mathbf{I}$
 where $\epsilon$ is a small scalar ($10^{-12}$) and $\mathbf{I}$ is the identity matrix, added to prevent $\boldsymbol{\Sigma}$ from being singular.
+Similarly, the asymptotic reinitialization is bypassed for cases where $N<D$.
+The effects of skipping the reinitialization can be seen in the $N=100$ sample-variant trials (Fig.
+4).
 New vectors $\mathbf{y}$ are generated by sampling from a multivariate normal distribution, with noise added to further promote exploration.
 (8) $\mathbf{y}\sim\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})+\boldsymbol{\delta}$
-The noise vector is $\boldsymbol{\delta}$ is sampled from a multivariate normal distribution, with respect to the search bounds:
-(9) $\boldsymbol{\delta}\sim\mathcal{N}\left(\mathbf{0},\left(\frac{\mathbf{b_{high,n}}-\mathbf{b_{low,n}}}{20}\right)^{2}\mathbf{I}\right)$
+The noise vector $\boldsymbol{\delta}$ is sampled from a multivariate normal distribution, with respect to the search bounds:
+(9) $\boldsymbol{\delta}\sim\mathcal{N}\left(0,\left(\frac{\mathbf{b}_{\text{high},n}-\mathbf{b}_{\text{low},n}}{20}\right)^{2}\mathbf{I}\right)$
 Finally, all components of the new solution $\mathbf{X}_{\text{new}}$ are clipped to ensure they remain within the defined search bounds $\mathbf{B}=[\mathbf{b}_{\text{low}},\mathbf{b}_{\text{high}}]$.
-(10) $\mathbf{X_{new,n}}=\text{clip}(\mathbf{y_{n}},\mathbf{B_{n}})$
+(10) $\mathbf{X}_{\text{new},n}=\text{clip}(\mathbf{y}_{n},\mathbf{B}_{n})$
 These new solutions completely replace the reinitialized solutions and are not subjected to crossover mechanisms.
-Due to the $\mathcal{O}(N)$ scaling of the covariance matrix calculations, along with decaying probability of occurrence, the additional overhead becomes negligible compared to the objective function evaluations.
+Due to the covariance calculations only using a fraction of the population, along with decaying probability of occurrence, the additional overhead becomes negligible compared to the objective function evaluations.
+For extreme cases, the reinitialization can be disabled, which is the default behavior for $N<D$ problems.
 Figure 2. QUASAR workflow for evolving each generation.
 3. Experimental Design
-3.1. Comparison Optimizers & Test Functions
+3.1. Optimizers & Test Functions
 The performance of QUASAR was benchmarked against modern implementations of Differential Evolution (DE) and L-SHADE, specifically SciPy’s DE and MealPy’s L-SHADE.
-This choice is justified by the fact that the QUASAR algorithm is built on the DE framework, as well as DE’s standing as a robust and effective optimizer (price1996differential).
-L-SHADE was specifically chosen due to its success in CEC optimization competitions (mo2019efficient).
+This choice is justified by the fact that the QUASAR algorithm is built on the DE framework, as well as DE’s standing as a robust and effective optimizer (Price, 1996).
+L-SHADE was specifically chosen due to its success in CEC optimization competitions (Mo et al., 2019).
 The trials were conducted using all 29 functions of the CEC2017 numerical optimization benchmark suite.
@@
 To ensure a fair comparison, each optimizer was run using their default hyperparameters.
-The initial population generation methods were also left to defaults; the tested sample sizes not being powers of 2 diminishes the uniformity of QUASAR’s initial Sobol sequence (joe2008constructing), providing a theoretical initial handicap.
-Since SciPy’s DE implementation internally scales the input population size by dimensionality ($popsize\times D$ when all dimensions have equal bounds), a correction factor of $1/D$ is applied to its initial sample size.
-The computational budget for metaheuristic comparison is often defined by the maximum number of objective function evaluations ($FES_{max}$) (kazikova2021objective).
+The initial population generation methods were also left to defaults; the tested sample sizes not being powers of 2 diminishes the uniformity of QUASAR’s initial Sobol sequence (Joe and Kuo, 2008), providing a theoretical initial handicap.
+Since SciPy’s DE implementation internally scales the input population size by dimensionality ($N\times D$ when all dimensions have equal bounds), a correction factor of $1/D$ is applied to its initial sample size.
+The computational budget for metaheuristic comparison is often defined by the maximum number of objective function evaluations ($FES_{max}$) (Kazikova et al., 2021).
 Due to L-SHADE’s linear population reduction, the $FES_{max}$ becomes a difficult comparison metric to compare against.
 As such, the primary comparisons are performed on final solution fitness errors and overall optimization times, given the same number of generations, using modern implementations via SciPy’s DE and MealPy’s L-SHADE.
-By default, the L-SHADE implementation features tolerance parameters for early convergence, and the QUASAR and DE implementations further include a final polishing step using L-BFGS-B minimization.
+By default, the L-SHADE implementation features tolerance parameters for early convergence.
+Additionally, the QUASAR and DE implementations include a final polishing step, using Powell and L-BFGS-B minimization, respectively.
 These were all disabled for the trials to ensure the optimizers fully utilize their computational budget, terminating only when reaching the maximum number of generations $g_{max}$.
@@
 The trials were structured into two distinct phases to assess QUASAR’s behavior under varying dimensionalities and population sizes.
-The results are compared based on the overall aggregated rank sums, geometric mean error factors ($GMERF$), and optimization run times.
-Dimension-Variant Trials:
-These trials assessed QUASAR’s scalability by varying the problem dimension $D$ with a fixed population size ($N=1000$).
-The dimension-variant experiments were run with $30$ independent trials for each function/dimension combination to ensure statistical significance.
+The results are compared based on the overall aggregated rank sums, geometric mean error factors ($R_{\text{error}}$), and optimization run time ratios $R_{\text{time}}$.
+3.3.1. Dimension-Variant Trials:
+The dimension-variant trials assessed QUASAR’s scalability by scaling the problem dimension $D$ with a fixed population size ($N=1000$).
+$30$ independent trials were performed for each test function/dimension combination to ensure statistical significance.
 The dimensions were varied between $[10,30,50,100]$, resulting in a total of 10,400 full optimization runs; 3,480 for each algorithm.
-The computational budget was kept constant, with maximum number of iterations $g_{max}=100$ and sample size $N=1000$.
-Table 1. Dimension-Variant Test Parameters
-Parameter Test Values
-Dimensions ($D$) [10, 30, 50, 100]
-Population Size ($N$) 1000
-Generations ($g_{\max}$) 100
-Sample-Variant Trials:
-These trials assessed the impact of population size on performance in high dimensions (a fixed $D=100$).
-The computational budget is varied between population sizes of [100, 250, 500, 1000, 5000], with 30 independent trials performed for each sample size / test function combination using maximum number of iterations $g_{max}=100$.
-The $N=5000$ trials, however, were only run for 10 trials to ensure statistical significance while mitigating the high computational overhead, resulting in a total of 11,310 full optimization runs; 3770 for each algorithm.
-Table 2. Sample-Variant Test Parameters
-Parameter Test Values
-Dimensions ($D$) 100
-Population Size ($N$) [100, 250, 500, 1000, 5000]
-Generations ($g_{\max}$) 100
+The computational budget was kept constant, with maximum number of iterations $g_{max}=100$ and fixed sample size $N=1000$.
+3.3.2. Sample-Variant Trials:
+The sample-variant trials assessed the impact of population size on performance.
+The computational budget is varied between population sizes of $[100,250,500,1000,5000]$.
+$30$ independent trials were performed for each test function / sample size combination, using maximum number of iterations $g_{max}=100$ and fixed dimensionality $D=100$.
+The $N=5000$ trials, however, were only run for 10 trials to ensure statistical significance while mitigating the high computational overhead, resulting in a total of 11,310 full optimization runs; 3770 for each optimization algorithm.
+Table 1. Experimental Parameters for Sample and Dimension Variations
+Sample-variant Dimension-variant
+D 100 [10, 30, 50, 100]
+N [100, 250, 500, 1000, 5000] 1000
+$\mathbf{g_{\max}}$ 100 100
 3.4. Statistical Analysis:
-Overall Rank:
+3.4.1. Overall Rank:
 To draw the overall statistical comparison between the optimizers across all function/dimension and function/sample size combinations, the Friedman test was used.
-For each unique combination, the median final error value ($f$) across all independent trials was calculated for each optimizer.
-Optimizers were ranked based on these median errors, with the best performance (lowest error) receiving a rank of 1.
+For each unique combination, the median final error value $f$ across all independent trials was calculated for each optimizer.
+The optimizers were ranked based on these median errors, with the best performance (lowest error) receiving a rank of 1.
 The ranks were summed across all scenarios to produce the aggregated rank sum for each optimizer. A lower sum indicates superior overall performance.
-Solution Quality Improvements:
-The primary comparison metric used to quantify the magnitude of performance difference is the Geometric Mean Error Reduction Factor $GMERF$, which provides an unbiased measure of relative performance across all $N_{\text{trials}}$ independent runs for a given scenario:
-$\text{GMERF}=\left(\prod_{i=1}^{N_{\text{trials}}}\frac{f_{\text{comparison},i}}{f_{\text{QUASAR},i}}\right)^{\frac{1}{N_{\text{trials}}}}$
-Here, $f_{\text{comparison},i}$ and $f_{\text{QUASAR},i}$ are the final error values achieved by the competitor algorithm (e.g., DE or L-SHADE) and QUASAR, respectively, in the $i$-th trial.
-GMERF>1{GMERF}>1 indicates that QUASAR performed better than the comparison optimizer.
-For the final aggregated results, the Overall $GMERF$ is calculated as the geometric mean of the GMERF values obtained for each function/dimension combination.
-$\text{GMERF}_{\text{Overall}}=\left(\prod_{j=1}^{N_{\text{combinations}}}\text{GMERF}_{j}\right)^{\frac{1}{N_{\text{combinations}}}}$
-Computational Efficiency:
+3.4.2. Solution Quality Improvements:
+The primary comparison metric used to quantify the magnitude of performance improvement is the geometric mean error reduction factor $R_{\text{error}}$.
+This provides an unbiased measure of relative final objective penalty $f$ for the comparison optimizer to QUASAR across all $T$ independent trials for each scenario:
+$R_{\text{error}}=\sqrt[T]{\prod_{i=1}^{T}\frac{f_{\text{comp},i}}{f_{\text{QUASAR},i}}}$
+Here, $f_{\text{comp},i}$ and $f_{\text{QUASAR},i}$ are the final error values achieved by the competitor algorithm (e.g., DE or L-SHADE) and QUASAR, respectively, in the $i$-th trial.
+For the overall results, the average $\bar{R}_{\text{error}}$ across all data is calculated as the geometric mean of the $R_{\text{error}}$ obtained for each of $k$ function/dimension or function/sample size combinations.
+${\bar{R}_{\text{error}}}=\sqrt[k]{\prod_{j=1}^{k}{R_{\text{error},j}}}$
+3.4.3. Computational Efficiency:
 The computational efficiency of the optimizers is quantified by comparing their average optimization run times.
-The ratios of run times $R_{t}$ for QUASAR vs.
-DE & L-SHADE are shown in Table 7 & Table 8, by dimension and sample size respectively.
-$R_{t}$ is calculated by dividing the average run time of the comparison method ($\text{Mean Sec}_{\text{comparison}}$) by the average run time of QUASAR ($\text{Mean Sec}_{\text{QUASAR}}$) across all trials for that specific dimension.
-(11) $R_{t}=\frac{\text{Mean Run Time}_{\text{comparison},D}}{\text{Mean Run Time}_{\text{QUASAR},D}}$
-The overall efficiency metrics are the overall run time ratios, $R_{t_{overall}}$, calculated as the arithmetic mean of each paired ratio $R_{t}$.
-(12) $\text{R}_{\text{t,overall}}=\frac{1}{|D|}\sum_{D\in D}\left(\frac{1}{|R_{t,D}|}\sum_{i=1}^{|R_{t,D}|}\frac{t_{\text{comparison},i,D}}{t_{\text{QUASAR},i,D}}\right)$
-Statistical Significance:
-For direct comparison within each dimension or sample size, the paired Wilcoxon signed-rank test is used.
-This non-parametric test was chosen as the final error distributions are on different scales for each function and are assumed not to be normal, and to ensure consistency with the non-parametric ranking methodology.
-The Wilcoxon test was conducted on the final error values ($f$) and run times to generate the p-values for each function/dimension combination.
+The run times ratios $R_{\text{time}}$ for QUASAR vs.
+DE & L-SHADE are shown in Table 5, by dimension and population size.
+$R_{\text{time}}$ is calculated by dividing the arithmetic average run time $\bar{t}$ of the comparison method by the average run time of QUASAR across all trials for that specific dimension.
+$R_{\text{time}}=\frac{\bar{t}_{\text{comp}}}{\bar{t}_{\text{QUASAR}}}=\frac{\sum_{i=1}^{T}t_{\text{comp},i}}{\sum_{i=1}^{T}t_{\text{QUASAR},i}}$
+The overall efficiency metric, $\bar{R}_{\text{time}}$, is defined as the geometric mean of the $R_{\text{time}}$ ratios across all $k$ function/dimension or function/sample size combinations.
+(11) $\bar{R}{\text{time}}=\sqrt[k]{\prod{j=1}^{k}R_{\text{time},j}}$
+3.4.4. Statistical Significance:
+For direct comparison within each dimension or sample size, the paired Wilcoxon signed-rank test is performed.
+This non-parametric test was chosen as the final error distributions are on different scales for each function, are assumed not to be normally distributed, and to ensure consistency with the non-parametric ranking methodology.
+The signed-rank test was conducted on the final error values ($f$) and run times to generate the p-values for each function/dimension combination.
 4. Results
-The primary solution quality metrics are the Friedman aggregated rank sum and the geometric mean error reduction factor $GMERF$, quantifying QUASAR’s improvements in final solution fitness.
-The computational efficiency is also measured by comparing full optimization run times.
-Since lower errors and run times indicate better performance, a $\text{ratio}>1$, for both the $GMERF$ and run times ratios, indicates that QUASAR demonstrated superior performance.
-Similarly, all plots shown are of raw fitness errors and run times; for all plots, lower values indicate higher performance.
+The primary solution quality metrics are the Friedman aggregated rank sum and the geometric mean error reduction factor $R_{\text{error}}$, quantifying QUASAR’s improvements in final solution fitness.
+The computational efficiency is measured by comparing the ratio of full optimization run times $R_{\text{time}}$.
+QUASAR demonstrated top performance in the rank sum test, shown in Table 2.
+Similarly, it demonstrated consistent improvements in geometric mean error reduction factor $R_{\text{error}}$ (Table 4) and run times $R_{\text{time}}$ (Table 5).
+Table 2. Rank Sums by Sample Size and Dimension
+N D Overall
+QUASAR 217 150 367
+L-SHADE 223 229 452
+DE 430 305 735
+Table 3. Overall Performance Comparison (Geometric Mean Ratios)
+vs L-SHADE vs DE
+$\mathbf{\bar{R}}$ 95% CI p-value $\mathbf{\bar{R}}$ 95% CI p-value
+Quality ($\mathbf{\bar{R}_{\text{error}}}$) 2.07 (0.62, 6.96) $5.30\times 10^{-144}$ 3.85 (0.93, 16.03) $5.78\times 10^{-136}$
+Run Time ($\mathbf{\bar{R}_{\text{time}}}$) 5.16 (4.57, 5.83) $3.36\times 10^{-141}$ 1.40 (1.36, 1.44) $4.52\times 10^{-23}$
+Ratios ¿ 1 indicate superior QUASAR performance.
 4.1. Solution Quality
-The ranks and geometric means of fitness error $GMERF$ of the final solution are calculated for both the dimension-variant and sample-variant trials.
-The 95% confidence intervals show consistent performance improvements in almost every scenario tested.
-4.1.1. Dimension-Variant Trials:
-The dimension-variant analysis assessed QUASAR’s convergence against DE and L-SHADE across four discrete dimensions ($10D$, $30D$, $50D$, and $100D$), using a fixed sample size $N=\text{1000}$ and maximum generations $g_{max}=100$.
-The geometric mean error reduction factors ($GMERF$) were consistently high across all dimensions (Fig 3).
-The 95% confidence intervals validate its consistent improvements; the intervals do not fall below $1.0\times$ GMERFGMERF.
-The largest improvements were seen against DE in $100D$, achieving a $GMERF$ of $\text{13.52}\times$ with a CI upper bound of $16.68\times$.
-The comparison against L-SHADE showed similar improvements; the full set of $GMERF$ values are shown in Table 5.
+The ranks and geometric mean fitness errors $R_{\text{error}}$ of the final solutions are calculated for both the dimension-variant and sample-variant trials.
+Solution quality comparisons are shown in Table 4.
+4.1.1. Dimension-Variant:
+The dimension-variant analysis assessed QUASAR’s convergence against DE and L-SHADE across four discrete dimensions $[10,30,50,100]$, using a fixed sample size $N=1000$ and maximum generations $g_{max}=100$.
+The geometric mean error reduction factors ($R_{\text{error}}$) were consistently high across all dimensions (Fig.
+3).
+The 95% confidence intervals validate its consistent improvements; the intervals do not fall below $R_{\text{error}}=1.0\times$.
+The largest improvements were seen against DE in $100D$, achieving a $R_{\text{error}}$ of $\text{13.52}\times$ with a CI upper bound of $16.68\times$.
+The comparison against L-SHADE showed the highest improvements in $30D$, with $R_{\text{error}}=4.14$, reaching CI upper bounds of 4.74.
 Figure 3. Geometric mean error (log-scale) by dimension for each optimizer, with population size of 1000.
-The Friedman test was used on the entire set of unique function/dimension combinations to determine the overall performance hierarchy of the optimizers.
-The test confirmed a statistically significant improvement in performance ($p\ll 0.001$), shown in Table 3.
-Table 3. Dimension-Variant Rank Sums
-Method Rank Sum
-QUASAR 150
-L-SHADE 229
-DE 305
-4.1.2. Sample Size-Variant Trials:
+4.1.2. Sample-Variant:
 To assess the impact of varying population sizes, QUASAR and DE were compared across a fixed $100D$ and 100 maximum generations $g_{max}$, using various sample sizes ($N$).
-The fitness improvement factors saw the highest gains with higher population sizes; in the $N=5000$ case, QUASAR reached a geometric mean error improvement factor $GMERF$ of $\text{18.5}\times$ compared to DE, and $4.47\times$ compared to L-SHADE.
-These performance increases scale super-linearly with population size (Fig.
-4, Table 6).
-The only scenario where QUASAR did not outperform L-SHADE was the $N=100$ (low-population) case, where L-SHADE achieved a $GMERF$ of $0.49\times$.
-The Friedman test is run on the set of all function/population size combinations, shown in Table 4.
+The fitness improvement factors saw the highest gains with higher population sizes; in the $N=5000$ case, QUASAR reached a geometric mean error improvement factor $R_{\text{error}}$ of $\text{18.5}\times$ compared to DE, and $4.47\times$ compared to L-SHADE.
+These performance increases scale with population size (Fig.
+4).
+The only scenario where QUASAR did not outperform L-SHADE was the $N=100$ case, where L-SHADE achieved $R_{\text{error}}=0.49\times$, highlighting the impact of bypassing the covariance reinitialization for $N<D$.
 Figure 4. Geometric mean error (log-scale) by population size for each optimizer, in 100 dimensions.
-Table 4. Sample-Variant Rank Sums
-Method Rank Sum
-QUASAR 217
-L-SHADE 223
-DE 430
-Table 5. Dimension-Variant GMERF Results: QUASAR vs L-SHADE and DE
-D vs L-SHADE vs DE p-value p-value QUASAR L-SHADE DE
-GMERF 95% CI GMERF 95% CI (vs L-SHADE) (vs DE) (GM Error) (GM Error) (GM Error)
-10 1.30 (1.20, 1.40) 1.06 (1.01, 1.12) $7\times 10^{-16}$ $7\times 10^{-3}$ $1.59\times 10^{3}$ $2.10\times 10^{3}$ $1.72\times 10^{3}$
-30 4.14 (3.62, 4.74) 4.94 (4.24, 5.77) $5\times 10^{-91}$ $3\times 10^{-108}$ $5.10\times 10^{3}$ $2.11\times 10^{4}$ $2.52\times 10^{4}$
-50 3.78 (3.32, 4.31) 8.44 (7.09, 10.04) $9\times 10^{-79}$ $5\times 10^{-134}$ $1.46\times 10^{4}$ $5.51\times 10^{4}$ $1.23\times 10^{5}$
-100 2.67 (2.32, 3.06) 13.52 (10.95, 16.68) $2\times 10^{-36}$ $2\times 10^{-141}$ $7.917\times 10^{4}$ $2.11\times 10^{5}$ $1.07\times 10^{6}$
-Table 6. Sample Size-Variant GMERF Results: QUASAR vs L-SHADE & DE
-S vs L-SHADE vs DE p-value p-value QUASAR L-SHADE DE
-GMERF 95% CI GMERF 95% CI (vs L-SHADE) (vs DE) (GM Error) (GM Error) (GM Error)
-100 0.49 (0.44, 0.56) 2.43 (2.04, 2.89) $7\times 10^{-86}$ $3\times 10^{-46}$ $5.94\times 10^{5}$ $2.94\times 10^{5}$ $1.44\times 10^{6}$
-250 1.34 (1.19, 1.50) 7.86 (6.51, 9.49) $3.3\times 10^{-2}$ $3\times 10^{-141}$ $1.79\times 10^{5}$ $2.39\times 10^{5}$ $1.41\times 10^{6}$
-500 1.93 (1.70, 2.20) 10.40 (8.54, 12.67) $2\times 10^{-8}$ $2\times 10^{-140}$ $1.15\times 10^{5}$ $2.21\times 10^{5}$ $1.19\times 10^{6}$
-1000 2.66 (2.32, 3.06) 13.52 (10.95, 16.68) $5\times 10^{-34}$ $2\times 10^{-141}$ $7.92\times 10^{4}$ $2.11\times 10^{5}$ $1.07\times 10^{6}$
-5000 4.47 (3.41, 5.86) 18.54 (12.49, 27.53) $3\times 10^{-37}$ $3\times 10^{-48}$ $4.20\times 10^{4}$ $1.88\times 10^{5}$ $7.78\times 10^{5}$
+Table 4. Solution Quality Comparison (Geometric Mean Error)
+vs L-SHADE vs DE
+$\mathbf{R_{error}}$ 95% CI p-value $\mathbf{R_{error}}$ 95% CI p-value
+$\mathbf{D=10}$ 1.30 (1.20, 1.40) $7.0\times 10^{-16}$ 1.06 (1.01, 1.12) $7.0\times 10^{-3}$
+$\mathbf{D=30}$ 4.14 (3.62, 4.74) $5.0\times 10^{-91}$ 4.94 (4.24, 5.77) $3.0\times 10^{-108}$
+$\mathbf{D=50}$ 3.78 (3.32, 4.31) $9.0\times 10^{-79}$ 8.44 (7.09, 10.04) $5.0\times 10^{-134}$
+$\mathbf{D=100}$ 2.67 (2.32, 3.06) $2.0\times 10^{-36}$ 13.52 (10.95, 16.68) $2.0\times 10^{-141}$
+$\mathbf{N=100}$ 0.49 (0.44, 0.56) $7.0\times 10^{-86}$ 2.43 (2.04, 2.89) $3.0\times 10^{-46}$
+$\mathbf{N=250}$ 1.34 (1.19, 1.50) $3.3\times 10^{-2}$ 7.86 (6.51, 9.49) $3.0\times 10^{-141}$
+$\mathbf{N=500}$ 1.93 (1.70, 2.20) $2.0\times 10^{-8}$ 10.40 (8.54, 12.67) $2.0\times 10^{-140}$
+$\mathbf{N=1000}$ 2.66 (2.32, 3.06) $5.0\times 10^{-34}$ 13.52 (10.95, 16.68) $2.0\times 10^{-141}$
+$\mathbf{N=5000}$ 4.47 (3.41, 5.86) $3.0\times 10^{-37}$ 18.54 (12.49, 27.53) $3.0\times 10^{-48}$
 4.2. Computational Efficiency
-The computational efficiency of QUASAR was measured using the overall ratio of run time $R_{t_{overall}}$ (seconds) relative to DE and L-SHADE.
-4.2.1. Dimension-Variant Trials:
-In the dimension-variant trials, QUASAR demonstrated consistently faster run times compared to DE and L-SHADE.
-The overall average run time improvement ratios $R_{t_{overall}}$ were $1.42\times$ and $6.29\times$ respectively.
-The average paired run time ratios $R_{t}$ are shown in Table 7, and optimization run times in seconds are in Table 7 and Fig.
-5.
+The computational efficiency of QUASAR was measured using the overall ratio of run time $R_{\text{time}}$ (seconds) relative to DE and L-SHADE. The full set of comparisons is shown in Table 5.
+4.2.1. Dimension-Variant:
+In the dimension-variant trials, QUASAR demonstrated consistently faster run times.
+The results, shown in Fig.
+5, consistently show that QUASAR achieved faster optimization speeds across all dimensions tested, scaling linearly with dimensionality.
 Figure 5. Average optimization run times for each optimizer, by dimension.
-4.2.2. Sample Size-Variant Trials:
-In the sample-variant trials, run times were consistently faster than DE and L-SHADE, averaging $7.78\times$ and $1.39\times$ speed increases respectively across all sample-variant trials (Fig.
-6, Table 8).
-The $N=250$ trials, however, demonstrated a ratio $R_{t}$ vs DE of $0.84\times$, indicating that DE was approximately 16% faster.
+4.2.2. Sample-Variant:
+In the sample-variant trials, run times were consistently faster than DE and L-SHADE, averaging $7.78\times$ and $1.39\times$ speed increases respectively across all sample-variant trials (Fig. 6).
+The $N=250$ trials, however, demonstrated a ratio $R_{\text{time}}$ vs.
+DE of $0.84\times$, indicating that DE was approximately 16% faster.
+This likely represents the overhead of the covariance matrix calculation relative to the low total evaluation count, whereas the covariance calculations are bypassed in the lower $N=100$ case where QUASAR was faster than DE.
 Figure 6. Average optimization run times for each optimizer, by population size.
-Table 7. Dimension-Variant Run Times: QUASAR vs L-SHADE & DE
-D $\mathbf{R_{t}}$ $\mathbf{R_{t}}$ p-value p-value QUASAR L-SHADE DE
-$\mathbf{vs\ \textbf{L-SHADE}}$ $\mathbf{vs\ DE}$ (vs L-SHADE) (vs DE) (sec) (sec) (sec)
-10 $4.21$ $1.26$ $5\text{\times}{10}^{-144}$ $7.3\text{\times}{10}^{-141}$ $16.4$ $69.1$ $20.7$
-30 $3.17$ $1.17$ $5\text{\times}{10}^{-144}$ $1.1\text{\times}{10}^{-139}$ $25.0$ $79.4$ $29.2$
-50 $2.66$ $1.14$ $5\text{\times}{10}^{-144}$ $1.6\text{\times}{10}^{-137}$ $33.5$ $89.0$ $38.1$
-100 $2.03$ $1.09$ $5\text{\times}{10}^{-144}$ $3\text{\times}{10}^{-137}$ $55.4$ $112.4$ $60.5$
-Table 8. Sample Size-Variant Run Times: QUASAR vs L-SHADE & DE
-S $\mathbf{R_{t}}$ $\mathbf{R_{t}}$ p-value p-value QUASAR L-SHADE DE
-$\mathbf{vs\ \textbf{L-SHADE}}$ $\mathbf{vs\ DE}$ (vs L-SHADE) (vs DE) (sec) (sec) (sec)
-100 $1.37$ $1.05$ $\text{\times}{10}^{-142} 1×1$$\text{\times}{10}^{-84}$ $5.8$ $7.9$ $6.0$
-250 $1.47$ $0.84$ $1\text{\times}{10}^{-142}$ $5\text{\times}{10}^{-58}$ $13.7$ $20.1$ $11.4$
-500 $1.65$ $1.06$ $1\text{\times}{10}^{-142}$ $2\text{\times}{10}^{-123}$ $28.7$ $47.2$ $30.5$
-1000 $2.00$ $1.08$ $1\text{\times}{10}^{-142}$ $9\text{\times}{10}^{-121}$ $57.3$ $114.4$ $62.0$
-5000 $5.96$ $1.29$ $3\text{\times}{10}^{-49}$ $3\text{\times}{10}^{-48}$ $292.0$ $1741.7$ $376.1$
+Table 5. Efficiency Comparison (Geometric Mean Run Time)
+vs L-SHADE vs DE GM Run Time (s)
+$\mathbf{R_{\text{time}}}$ p-value $\mathbf{R_{\text{time}}}$ p-value QUASAR L-SHADE DE
+$\mathbf{D=10}$ 4.21 $5\times 10^{-144}$ 1.26 $7.3\times 10^{-141}$ 16.4 69.1 20.7
+$\mathbf{D=30}$ 3.17 $5\times 10^{-144}$ 1.17 $1.1\times 10^{-139}$ 25.0 79.4 29.2
+$\mathbf{D=50}$ 2.66 $5\times 10^{-144}$ 1.14 $1.6\times 10^{-137}$ 33.5 89.0 38.1
+$\mathbf{D=100}$ 2.03 $5\times 10^{-144}$ 1.09 $3\times 10^{-137}$ 55.4 112.4 60.5
+$\mathbf{N=100}$ 1.37 $1\times 10^{-142}$ 1.05 $1\times 10^{-84}$ 5.8 7.9 6.0
+$\mathbf{N=250}$ 1.47 $1\times 10^{-142}$ 0.84 $5\times 10^{-58}$ 13.7 20.1 11.4
+$\mathbf{N=500}$ 1.65 $1\times 10^{-142}$ 1.06 $2\times 10^{-123}$ 28.7 47.2 30.5
+$\mathbf{N=1000}$ 2.00 $1\times 10^{-142}$ 1.08 $9\times 10^{-121}$ 57.3 114.4 62.0
+$\mathbf{N=5000}$ 5.96 $3\times 10^{-49}$ 1.29 $3\times 10^{-48}$ 292.0 1741.7 376.1
 5. Ease of Use and Implementation
-QUASAR is designed to be robust with minimal parameter dependence, making it effective for most numerical optimization problems.
 5.1. Ease of Use
-A key feature of QUASAR is its ability to operate effectively with its default entanglerateentangle\ rate, bypassing the hyperparameter tuning often required by other metaheuristics such as DE and L-SHADE.
+QUASAR operates most effectively with its default parameters, bypassing the hyperparameter tuning often required by other metaheuristics.
 The algorithm relies on its quasi-adaptive mechanisms to dynamically balance exploration and exploitation.
-For typical use cases, the only inputs needed are the objective function, search bounds, and the target number of generations ($g_{\max}$).
+Typical use cases require only the objective function, search bounds, and the target number of generations ($g_{\max}$).
+For additional control, the entangle rate can be tuned between $[0,1]$, though the default value of 0.33 was most robust across the CEC benchmarks.
+While not used in this study, the optimizer naturally handles vectorized objective functions by evolving the full population at once.
+This can lead to orders of magnitude faster run times, depending on population size.
 5.2. Implementation
-QUASAR is implemented as part of a streamlined open-source Python package, designed for seamless integration into existing workflows.
-The syntax follows the structure of SciPy’s optimization module, making it nearly one-to-one interchangeable with its Differential Evolution.
+QUASAR is implemented as part of a streamlined open-source Python package, designed for ease of use and adoption into existing workflows.
+The syntax follows the structure of SciPy’s optimize module, making it nearly one-to-one interchangeable with its Differential Evolution function.
 The source code requires minimal dependencies, relying only on NumPy and SciPy.
-The open-source package containing QUASAR, hdim-opt, is available via the Python Package Index (PyPI) using ’pip install hdim_opt’, or at github.com/jgsoltes/hdim-opt which includes the experimental notebooks for full reproducibility.
+The open-source package containing QUASAR, hdim_opt, is available via PyPi or at https://github.com/jgsoltes/hdim-opt, which includes the experimental notebooks for full reproducibility.
 6. Conclusion
-This study introduced Quasi-Adaptive Search with Asymptotic Reinitialization (QUASAR), a quantum-inspired evolutionary algorithm designed to efficiently explore and exploit high-dimensional spaces.
-The algorithm utilizes a highly stochastic architecture, including probabilistic mutation mechanisms, rank-based crossover rates, and asymptotically decaying covariance reinitializations.
-Experimental trials across the CEC2017 benchmark suite confirm that QUASAR delivers exceptional performance compared to Differential Evolution (DE) and L-SHADE, consistently ranking best via the Friedman test.
-[…]
-The computational efficiency of the algorithm is a key benefit, with run times consistently lower than the comparison optimizers ($R_{t}>1$), scaling positively with population size and dimensionality.
+This study introduced Quasi-Adaptive Search with Asymptotic Reinitialization (QUASAR), a quantum-inspired evolutionary algorithm designed to efficiently explore and exploit complex high-dimensional spaces.
+Combining probabilistic mutation, rank-based crossover, and asymptotically decaying covariance reinitializations, QUASAR’s highly stochastic architecture is exceptionally suited for discontinuous and non-parametric optimization.
+Its reliance on default parameters and streamlined implementation allows for seamless integration within existing workflows.
+Experimental trials across the CEC2017 benchmark suite confirm that QUASAR delivers exceptional performance, achieving the highest rank via the Friedman test compared to Differential Evolution (DE) and L-SHADE, with respective overall geometric mean error improvements of $3.85\times$ and $2.07\times$.
+QUASAR also proves computationally efficient, with overall geometric mean run time improvements of $1.40\times$ and $5.16\times$ over DE and L-SHADE.
+Analysis by dimension and population size shows linear improvement scaling across all metrics.
 The empirical evidence validates QUASAR as a powerful evolutionary algorithm, ready for the next generation of high-dimensional optimization problems.
 7. Future Works
-7.1. Enhancements:
+7.1. Enhancements
 QUASAR’s quasi-adaptive mechanisms are highly effective; future improvements will explore fully self-adaptive strategies, similar to those found in other modern DE variants.
-The current covariance matrix ($\mathbf{\Sigma}$) treats all top solutions equall;.
-future work will investigate applying weights to the very best solutions.
+7.1.1. Mutations:
+The mutation factors are sampled from fixed Gaussian distributions; dynamic scaling based on convergence may significantly help to maintain a healthy balance of exploration and exploitation.
+The strategies may benefit from different permutations of equation terms, or non-probablistic selction methods.
+7.1.2. Crossover:
+Crossover rate is currently defined by fitness rank, and may benefit from a probability-based approach similar to the other QUASAR components, via soft-max or similar methods.
+Many techniques from existing evolutionary frameworks may prove beneficial to the model.
+7.1.3. Reinitialization:
+The current reinitialization covariance matrix ($\mathbf{\Sigma}$) treats all top solutions equally; future work will investigate weighted solutions similar to those found in CMA-ES.
 New methods for defining the probability of occurrence may be explored, such as sinusoidal or sigmoid activation behavior.
-Crossover rate is currently defined by fitness rank, and may benefit from a probability-based approach, similar to the other QUASAR components, via soft-max or similar methods.
-7.2. Verification:
-QUASAR’s highly stochastic framework makes it a practical choice for non-differentiable and non-parametric problems.
-The algorithms’s robustness would be further validated by rigorously evaluating its performance on more constrained, non-parametric problems.
+Additionally, the reinitialization is currently bypassed for problems where $N<D$. Testing other redistribution methods, such as uniform quasi-Monte Carlo sequences, may be beneficial for these cases.
+7.1.4. Verification:
+While QUASAR’s highly stochastic framework makes it a practical choice for non-differentiable and non-parametric problems, the algorithms’s performance would be further validated by rigorously evaluating its performance on constrained or multi-objective problems.
 Acknowledgements.

### v4

@@
-An Evolutionary Algorithm to Accelerate High-Dimensional Optimization
-Abstract.
+An Evolutionary Algorithm to Accelerate High-Dimensional Numerical Optimization
+Abstract
 High-dimensional numerical optimization presents a persistent challenge in computational science.
@@
 QUASAR’s performance establishes it as an effective, efficient, and user-friendly evolutionary algorithm for complex high-dimensional problems.
-copyright: none
-1. Introduction
-High-dimensional numerical optimization remains a persistent challenge across many fields, including machine learning, science, engineering, and finance (Zhou et al., 2024).
-This challenge stems from the curse of dimensionality; the exponential growth of the search space renders gradient-based and parametric search methods ineffective (Shan and Wang, 2008).
+keywords:
+Evolutionary computation , Metaheuristics , Global optimization , Quantum-inspired algorithm , Covariance reinitialization , High-dimensional optimization
+journal: Swarm and Evolutionary Computation
+[…]
+organization=Regis University,addressline=3333 Regis Blvd., city=Denver, postcode=80221, state=Colorado, country=USA
+1 Introduction
+High-dimensional numerical optimization remains a persistent challenge across many fields, including machine learning, science, engineering, and finance [9].
+This challenge stems from the curse of dimensionality; the exponential growth of the search space renders gradient-based and parametric search methods ineffective [6].
 Consequently, optimization metaheuristics, particularly evolutionary algorithms, have become essential for locating optimal solutions in complex function landscapes.
-Among these methods, Differential Evolution (DE) has established itself as a robust and highly effective optimizer due to its simplicity, efficiency, and effective mutation strategies (Storn, 1996).
+Among these methods, Differential Evolution (DE) has established itself as a robust and highly effective optimizer due to its simplicity, efficiency, and effective mutation strategies [8].
 However, standard DE variants can struggle to maintain an effective balance between global exploration and local exploitation, particularly in high-dimensional search spaces.
@@
 The algorithm’s primary innovation is an asymptotically decaying reinitialization mechanism, utilizing the spatial covariance matrix of top-performing solutions to inject high-quality genetic diversity into the population.
-The efficacy of QUASAR is demonstrated through rigorous testing against standard DE and its state-of-the-art variant L-SHADE across the complete CEC2017 benchmark suite.
-The results show large and statistically significant performance gains; these improvements are achieved using default parameters, highlighting QUASAR’s ease of use.
-The remainder of this paper is organized as follows: Section 2 details the QUASAR algorithm, including its mutation, crossover, selection, and reinitialization mechanisms.
-Section 3 defines the experimental design, and Section 4 analyzes the experimental results compared to DE and L-SHADE.
+The efficacy of QUASAR is demonstrated through rigorous evaluation against standard DE and its state-of-the-art variant L-SHADE across the full CEC2017 test function benchmark suite.
+The results show large and statistically significant improvements in final solution quality, with high computational efficiency.
+These performance gains are seen using its default parameters, highlighting the ease of use and scalability of the algorithm.
+The remainder of the paper is structured as follows: Section 2 details the QUASAR algorithm, including its mutation, crossover, selection, and reinitialization mechanisms.
+Section 3 defines the experimental design and statistical analysis, Section 4 analyzes the results.
 Section 5 describes the ease of use and implementation of the algorithm.
-Section 6 concludes the paper, with Section 7 proposing future work.
-2. QUASAR Algorithm
+Section 6 summarizes the results and concludes the paper, with Section 7 proposing future enhancements and validations.
+2 QUASAR Algorithm
 Quasi-Adaptive Search with Asymptotic Reinitialization (QUASAR) operates using a highly stochastic, multi-strategy approach.
 The mechanisms defined below synergize to create a dynamic ’push-and-pull’ effect on the solution vectors.
-2.1. Initial Population
-The initial population is generated as a Sobol sequence by default, due to its low discrepancy and superior uniformity in high-dimensional spaces compared to other quasi-Monte Carlo sampling methods (Bonneel et al., 2025).
+2.1 Initial Population
+The initial population is generated as a Sobol sequence by default, due to its low discrepancy and superior uniformity in high-dimensional spaces compared to other quasi-Monte Carlo sampling methods [1].
 The default population size, $N$, is set to $10\times D$, where $D$ is the number of dimensions.
-Additional initialization methods include uniform Latin Hypercube and random distributions, as well as non-uniform Hyperellipsoid Density sequences (Soltes, 2025).
-2.2. Mutation
+Additional initialization methods include uniform Latin Hypercube and random distributions, as well as non-uniform Hyperellipsoid Density sequences [7].
+2.2 Mutation
 Solutions are recombined with one another, analogous to particles sharing information via quantum entanglement.
@@
 This value defaults to $0.33$, resulting in the three strategies being applied equally across the population.
-2.2.1. Local Mutation (Exploitation):
+2.2.1 Local Mutation (Exploitation):
 The local mutation strategy, ’Spooky-Best’, is chosen with probability entangle rate.
 This strategy is similar to the DE/best/1 format, exploiting the immediate search space around the current best solution $\mathbf{X}_{\text{best},g}$.
-(Spooky-Best) $\mathbf{v}_{i,g}=\mathbf{X}_{\text{best},g}+F_{\text{local}}(\mathbf{X}_{i,g}-\mathbf{X}_{\text{rand},g})$
+$\mathbf{v}_{i,g}=\mathbf{X}_{\text{best},g}+F_{\text{local}}(\mathbf{X}_{i,g}-\mathbf{X}_{\text{rand},g})$ (Spooky-Best)
 The associated mutation factor $F_{\text{local}}$ is sampled from a normal distribution $\mathcal{N}(0,0.33^{2})$, allowing for small bi-directional perturbations about the best solution.
-2.2.2. Global Mutations (Exploration):
+2.2.2 Global Mutations (Exploration):
 Two global mutation strategies are applied to the remaining $(1-\textit{entangle rate})$ fraction of the population.
 Promoting global exploration, these strategies, ’Spooky-Current’ and ’Spooky-Random’, each have a 50% probability of being selected:
-(Spooky-Current) $\mathbf{v}_{i,g}=\mathbf{X}_{i,g}+F_{\text{global}}(\mathbf{X}_{\text{best},g}-\mathbf{X}_{\text{rand},g})$
-(Spooky-Random) $\mathbf{v}_{i,g}=\mathbf{X}_{\text{rand},g}+F_{\text{global}}(\mathbf{X}_{i,g}-\mathbf{X}_{\text{rand},g})$
+$\mathbf{v}_{i,g}=\mathbf{X}_{i,g}+F_{\text{global}}(\mathbf{X}_{\text{best},g}-\mathbf{X}_{\text{rand},g})$ (Spooky-Current)
+$\mathbf{v}_{i,g}=\mathbf{X}_{\text{rand},g}+F_{\text{global}}(\mathbf{X}_{i,g}-\mathbf{X}_{\text{rand},g})$ (Spooky-Random)
 The mutation factor $F_{\text{global}}$ is sampled from a bimodal distribution $\mathcal{N}(0.5,0.25^{2})+\mathcal{N}(-0.5,0.25^{2})$ to drive non-local exploration.
@@
 The terms used in the three mutation strategies are defined as:
+  • 1.
     $\mathbf{v}_{i,g}$: Mutant vector for the $i$-th individual at generation $g$.
+  • 2.
     $\mathbf{X}_{i,g}$: Current solution vector.
+  • 3.
     $\mathbf{X}_{\text{best},g}$: Best known solution in the population.
+  • 4.
     $\mathbf{X}_{\text{rand},g}$: Randomly chosen solution vector.
+  • 5.
     $F_{\text{local}}$ and $F_{\text{global}}$: Corresponding mutation factors.
-Figure 1. Mutation factor distributions for global and local mutation strategies.
-2.3. Crossover
-QUASAR uses binomial crossover with a dynamic crossover rate ($CR_{i}$) proportional to the solution’s fitness rank.
-The best solutions are rapidly recombined to maximize both exploration and exploitation via synergy with global and local mutations.
+Figure 1: Mutation factor distributions for global and local mutation strategies.
+2.3 Crossover
+QUASAR uses binomial crossover with a dynamic crossover rate ($CR_{i}$) inversely proportional to the solution’s fitness rank.
+The best solutions are rapidly recombined to maximize both exploration and exploitation through synergy with global and local mutations.
 The raw crossover rate ($CR_{\text{raw}},i$) for the $i$-th solution is calculated based on its rank ($\textit{rank}_{i}\in[0,N-1]$), where $\textit{rank}=0$ corresponds to the best solution (lowest objective function value).
@@
 This mechanism ensures that the best solutions have a higher chance of inheriting components from the new mutant vector $\mathbf{v}_{i}$.
-(1) $CR_{\text{raw},i}=\frac{N-1-\text{rank}_{i}}{N-1}$
-(2) $CR_{i}=\max(CR_{\text{raw},i},0.33)$
+$CR_{\text{raw},i}=\frac{N-1-\text{rank}_{i}}{N-1}$ (1)
+$CR_{i}=\max(CR_{\text{raw},i},0.33)$ (2)
 The trial vector $\mathbf{u}_{i}$ is then generated component-wise ($n$ dimension) using binomial crossover, comparing a uniform random number $\text{rand}(0,1)$ against $CR_{i}$:
-(3) $u_{i,n}=\begin{cases}v_{i,n},&\text{if }\text{rand}(0,1)\leq CR_{i}\\ X_{i,n},&\text{if }\text{rand}(0,1)>CR_{i}\end{cases}$
-2.4. Selection
+$u_{i,n}=\begin{cases}v_{i,n},&\text{if }\text{rand}(0,1)\leq CR_{i}\\ X_{i,n},&\text{if }\text{rand}(0,1)>CR_{i}\end{cases}$ (3)
+2.4 Selection
 A greedy elitism strategy is used to determine the next generation’s population, where the target vector for generation $g+1$ is the better of the two solutions between the current vector $\mathbf{X}_{i,g}$ and the trial vector $\mathbf{u}_{i,g}$.
 If a solution undergoes reinitialization, it bypasses the selection step.
-(4) $\mathbf{X}_{i,(g+1)}=\begin{cases}\mathbf{u}_{i},&\text{if }f(\mathbf{u}_{i})<f(\mathbf{X}_{i})\\ \mathbf{X}_{i},&\text{if }f(\mathbf{u}_{i})\geq f(\mathbf{X}_{i})\end{cases}$
-2.5. Asymptotic Reinitialization
+$\mathbf{X}_{i,(g+1)}=\begin{cases}\mathbf{u}_{i},&\text{if }f(\mathbf{u}_{i})<f(\mathbf{X}_{i})\\ \mathbf{X}_{i},&\text{if }f(\mathbf{u}_{i})\geq f(\mathbf{X}_{i})\end{cases}$ (4)
+2.5 Asymptotic Reinitialization
 The covariance-guided vector reinitialization applies to a fixed 33% of the worst-performing solutions.
@@
 The probability of reinitialization $P_{\text{reinit}}(g)$ decays asymptotically from $P_{\text{reinit}}(0)=1$ toward a target final probability, calculated using the current generation $g$:
-(5) $P_{\text{reinit}}(g)=\exp\left(\frac{\ln(P_{\text{final}})}{g_{\text{final}}\cdot g_{\text{max}}}\cdot g\right)$
+$P_{\text{reinit}}(g)=\exp\left(\frac{\ln(P_{\text{final}})}{g_{\text{final}}\cdot g_{\text{max}}}\cdot g\right)$ (5)
 where $g_{\text{max}}$ is the maximum generation count, $P_{\text{final}}=0.33$ is the target final probability, and $g_{\text{final}}=0.33$ defines the fraction of generations at which $P_{reinit}$ reaches $P_{final}$.
@@
 The top $M$ solutions (where $M=25\%\text{of the population}$) are used to calculate the mean vector $\boldsymbol{\mu}$ and the covariance matrix $\boldsymbol{\Sigma}$:
-(6) $\boldsymbol{\mu}=\frac{1}{M}\sum_{i=1}^{M}\mathbf{X}_{i}$
-(7) $\boldsymbol{\Sigma}=\frac{1}{M-1}\sum_{i=1}^{M}(\mathbf{X}_{i}-\boldsymbol{\mu})(\mathbf{X}_{i}-\boldsymbol{\mu})^{\top}+\epsilon\mathbf{I}$
+$\boldsymbol{\mu}=\frac{1}{M}\sum_{i=1}^{M}\mathbf{X}_{i}$ (6)
+$\boldsymbol{\Sigma}=\frac{1}{M-1}\sum_{i=1}^{M}(\mathbf{X}_{i}-\boldsymbol{\mu})(\mathbf{X}_{i}-\boldsymbol{\mu})^{\top}+\epsilon\mathbf{I}$ (7)
 where $\epsilon$ is a small scalar ($10^{-12}$) and $\mathbf{I}$ is the identity matrix, added to prevent $\boldsymbol{\Sigma}$ from being singular.
@@
 New vectors $\mathbf{y}$ are generated by sampling from a multivariate normal distribution, with noise added to further promote exploration.
-(8) $\mathbf{y}\sim\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})+\boldsymbol{\delta}$
+$\mathbf{y}\sim\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})+\boldsymbol{\delta}$ (8)
 The noise vector $\boldsymbol{\delta}$ is sampled from a multivariate normal distribution, with respect to the search bounds:
-(9) $\boldsymbol{\delta}\sim\mathcal{N}\left(0,\left(\frac{\mathbf{b}_{\text{high},n}-\mathbf{b}_{\text{low},n}}{20}\right)^{2}\mathbf{I}\right)$
+$\boldsymbol{\delta}\sim\mathcal{N}\left(0,\left(\frac{\mathbf{b}_{\text{high},n}-\mathbf{b}_{\text{low},n}}{20}\right)^{2}\mathbf{I}\right)$ (9)
 Finally, all components of the new solution $\mathbf{X}_{\text{new}}$ are clipped to ensure they remain within the defined search bounds $\mathbf{B}=[\mathbf{b}_{\text{low}},\mathbf{b}_{\text{high}}]$.
-(10) $\mathbf{X}_{\text{new},n}=\text{clip}(\mathbf{y}_{n},\mathbf{B}_{n})$
+$\mathbf{X}_{\text{new},n}=\text{clip}(\mathbf{y}_{n},\mathbf{B}_{n})$ (10)
 These new solutions completely replace the reinitialized solutions and are not subjected to crossover mechanisms.
@@
 For extreme cases, the reinitialization can be disabled, which is the default behavior for $N<D$ problems.
-Figure 2. QUASAR workflow for evolving each generation.
-3. Experimental Design
-3.1. Optimizers & Test Functions
+Figure 2: QUASAR workflow for evolving each generation.
+3 Experimental Design
+3.1 Optimizers & Test Functions
 The performance of QUASAR was benchmarked against modern implementations of Differential Evolution (DE) and L-SHADE, specifically SciPy’s DE and MealPy’s L-SHADE.
-This choice is justified by the fact that the QUASAR algorithm is built on the DE framework, as well as DE’s standing as a robust and effective optimizer (Price, 1996).
-L-SHADE was specifically chosen due to its success in CEC optimization competitions (Mo et al., 2019).
+This choice is justified by the fact that the QUASAR algorithm is built on the DE framework, as well as DE’s standing as a robust and effective optimizer [5].
+L-SHADE was specifically chosen due to its success in CEC optimization competitions [4].
 The trials were conducted using all 29 functions of the CEC2017 numerical optimization benchmark suite.
-3.2. Experimental Setup
+3.2 Experimental Setup
 To ensure a fair comparison, each optimizer was run using their default hyperparameters.
-The initial population generation methods were also left to defaults; the tested sample sizes not being powers of 2 diminishes the uniformity of QUASAR’s initial Sobol sequence (Joe and Kuo, 2008), providing a theoretical initial handicap.
+The initial population generation methods were also left to defaults; the tested sample sizes not being powers of 2 diminishes the uniformity of QUASAR’s initial Sobol sequence [2], providing a theoretical initial handicap.
 Since SciPy’s DE implementation internally scales the input population size by dimensionality ($N\times D$ when all dimensions have equal bounds), a correction factor of $1/D$ is applied to its initial sample size.
-The computational budget for metaheuristic comparison is often defined by the maximum number of objective function evaluations ($FES_{max}$) (Kazikova et al., 2021).
+The computational budget for metaheuristic comparison is often defined by the maximum number of objective function evaluations ($FES_{max}$) [3].
 Due to L-SHADE’s linear population reduction, the $FES_{max}$ becomes a difficult comparison metric to compare against.
@@
 These were all disabled for the trials to ensure the optimizers fully utilize their computational budget, terminating only when reaching the maximum number of generations $g_{max}$.
-3.3. Experimental Trials:
+3.3 Experimental Trials:
 The trials were structured into two distinct phases to assess QUASAR’s behavior under varying dimensionalities and population sizes.
 The results are compared based on the overall aggregated rank sums, geometric mean error factors ($R_{\text{error}}$), and optimization run time ratios $R_{\text{time}}$.
-3.3.1. Dimension-Variant Trials:
+3.3.1 Dimension-Variant Trials:
 The dimension-variant trials assessed QUASAR’s scalability by scaling the problem dimension $D$ with a fixed population size ($N=1000$).
-$30$ independent trials were performed for each test function/dimension combination to ensure statistical significance.
+3030 independent trials were performed for each test function/dimension combination to ensure statistical significance.
 The dimensions were varied between $[10,30,50,100]$, resulting in a total of 10,400 full optimization runs; 3,480 for each algorithm.
 The computational budget was kept constant, with maximum number of iterations $g_{max}=100$ and fixed sample size $N=1000$.
-3.3.2. Sample-Variant Trials:
+3.3.2 Sample-Variant Trials:
 The sample-variant trials assessed the impact of population size on performance.
 The computational budget is varied between population sizes of $[100,250,500,1000,5000]$.
-$30$ independent trials were performed for each test function / sample size combination, using maximum number of iterations $g_{max}=100$ and fixed dimensionality $D=100$.
+3030 independent trials were performed for each test function / sample size combination, using maximum number of iterations $g_{max}=100$ and fixed dimensionality $D=100$.
 The $N=5000$ trials, however, were only run for 10 trials to ensure statistical significance while mitigating the high computational overhead, resulting in a total of 11,310 full optimization runs; 3770 for each optimization algorithm.
-Table 1. Experimental Parameters for Sample and Dimension Variations
+Table 1: Experimental Parameters for Sample and Dimension Variations
 Sample-variant Dimension-variant
@@
 $\mathbf{g_{\max}}$ 100 100
-3.4. Statistical Analysis:
-3.4.1. Overall Rank:
+3.4 Statistical Analysis:
+3.4.1 Overall Rank:
 To draw the overall statistical comparison between the optimizers across all function/dimension and function/sample size combinations, the Friedman test was used.
@@
 The ranks were summed across all scenarios to produce the aggregated rank sum for each optimizer. A lower sum indicates superior overall performance.
-3.4.2. Solution Quality Improvements:
+3.4.2 Solution Quality Improvements:
 The primary comparison metric used to quantify the magnitude of performance improvement is the geometric mean error reduction factor $R_{\text{error}}$.
@@
 ${\bar{R}_{\text{error}}}=\sqrt[k]{\prod_{j=1}^{k}{R_{\text{error},j}}}$
-3.4.3. Computational Efficiency:
+3.4.3 Computational Efficiency:
 The computational efficiency of the optimizers is quantified by comparing their average optimization run times.
@@
 The overall efficiency metric, $\bar{R}_{\text{time}}$, is defined as the geometric mean of the $R_{\text{time}}$ ratios across all $k$ function/dimension or function/sample size combinations.
-(11) $\bar{R}{\text{time}}=\sqrt[k]{\prod{j=1}^{k}R_{\text{time},j}}$
-3.4.4. Statistical Significance:
+$\bar{R}_{\text{time}}=\sqrt[k]{\prod{j=1}^{k}R_{\text{time},j}}$ (11)
+3.4.4 Statistical Significance:
 For direct comparison within each dimension or sample size, the paired Wilcoxon signed-rank test is performed.
@@
 The signed-rank test was conducted on the final error values ($f$) and run times to generate the p-values for each function/dimension combination.
-4. Results
+4 Results
 The primary solution quality metrics are the Friedman aggregated rank sum and the geometric mean error reduction factor $R_{\text{error}}$, quantifying QUASAR’s improvements in final solution fitness.
@@
 Similarly, it demonstrated consistent improvements in geometric mean error reduction factor $R_{\text{error}}$ (Table 4) and run times $R_{\text{time}}$ (Table 5).
-Table 2. Rank Sums by Sample Size and Dimension
+Table 2: Rank Sums by Sample Size and Dimension
 N D Overall
@@
 DE 430 305 735
-Table 3. Overall Performance Comparison (Geometric Mean Ratios)
+Table 3: Overall Performance Comparison (Geometric Mean Ratios)
 vs L-SHADE vs DE
@@
 Ratios ¿ 1 indicate superior QUASAR performance.
-4.1. Solution Quality
+4.1 Solution Quality
 The ranks and geometric mean fitness errors $R_{\text{error}}$ of the final solutions are calculated for both the dimension-variant and sample-variant trials.
 Solution quality comparisons are shown in Table 4.
-4.1.1. Dimension-Variant:
+4.1.1 Dimension-Variant:
 The dimension-variant analysis assessed QUASAR’s convergence against DE and L-SHADE across four discrete dimensions $[10,30,50,100]$, using a fixed sample size $N=1000$ and maximum generations $g_{max}=100$.
@@
 The comparison against L-SHADE showed the highest improvements in $30D$, with $R_{\text{error}}=4.14$, reaching CI upper bounds of 4.74.
-Figure 3. Geometric mean error (log-scale) by dimension for each optimizer, with population size of 1000.
-4.1.2. Sample-Variant:
-To assess the impact of varying population sizes, QUASAR and DE were compared across a fixed $100D$ and 100 maximum generations $g_{max}$, using various sample sizes ($N$).
+Figure 3: Geometric mean error (log-scale) by dimension for each optimizer, with population size of 1000.
+4.1.2 Sample-Variant:
+To assess the impact of varying population sizes, the optimizers were compared across a fixed $100D$ and 100 maximum generations $g_{max}$, using various sample sizes ($N$).
 The fitness improvement factors saw the highest gains with higher population sizes; in the $N=5000$ case, QUASAR reached a geometric mean error improvement factor $R_{\text{error}}$ of $\text{18.5}\times$ compared to DE, and $4.47\times$ compared to L-SHADE.
@@
 The only scenario where QUASAR did not outperform L-SHADE was the $N=100$ case, where L-SHADE achieved $R_{\text{error}}=0.49\times$, highlighting the impact of bypassing the covariance reinitialization for $N<D$.
-Figure 4. Geometric mean error (log-scale) by population size for each optimizer, in 100 dimensions.
-Table 4. Solution Quality Comparison (Geometric Mean Error)
+Figure 4: Geometric mean error (log-scale) by population size for each optimizer, in 100 dimensions.
+Table 4: Solution Quality Comparison (Geometric Mean Error)
 vs L-SHADE vs DE
@@
 $\mathbf{N=5000}$ 4.47 (3.41, 5.86) $3.0\times 10^{-37}$ 18.54 (12.49, 27.53) $3.0\times 10^{-48}$
-4.2. Computational Efficiency
+4.2 Computational Efficiency
 The computational efficiency of QUASAR was measured using the overall ratio of run time $R_{\text{time}}$ (seconds) relative to DE and L-SHADE. The full set of comparisons is shown in Table 5.
-4.2.1. Dimension-Variant:
+4.2.1 Dimension-Variant Efficiency:
 In the dimension-variant trials, QUASAR demonstrated consistently faster run times.
@@
 5, consistently show that QUASAR achieved faster optimization speeds across all dimensions tested, scaling linearly with dimensionality.
-Figure 5. Average optimization run times for each optimizer, by dimension.
-4.2.2. Sample-Variant:
+Figure 5: Average optimization run times for each optimizer, by dimension.
+4.2.2 Sample-Variant Efficiency:
 In the sample-variant trials, run times were consistently faster than DE and L-SHADE, averaging $7.78\times$ and $1.39\times$ speed increases respectively across all sample-variant trials (Fig. 6).
@@
 This likely represents the overhead of the covariance matrix calculation relative to the low total evaluation count, whereas the covariance calculations are bypassed in the lower $N=100$ case where QUASAR was faster than DE.
-Figure 6. Average optimization run times for each optimizer, by population size.
-Table 5. Efficiency Comparison (Geometric Mean Run Time)
+Figure 6: Average optimization run times for each optimizer, by population size.
+Table 5: Efficiency Comparison (Geometric Mean Run Time)
 vs L-SHADE vs DE GM Run Time (s)
@@
 $\mathbf{N=5000}$ 5.96 $3\times 10^{-49}$ 1.29 $3\times 10^{-48}$ 292.0 1741.7 376.1
-5. Ease of Use and Implementation
-5.1. Ease of Use
+5 Ease of Use and Implementation
+5.1 Ease of Use
 QUASAR operates most effectively with its default parameters, bypassing the hyperparameter tuning often required by other metaheuristics.
@@
 This can lead to orders of magnitude faster run times, depending on population size.
-5.2. Implementation
+5.2 Implementation
 QUASAR is implemented as part of a streamlined open-source Python package, designed for ease of use and adoption into existing workflows.
@@
 The source code requires minimal dependencies, relying only on NumPy and SciPy.
-The open-source package containing QUASAR, hdim_opt, is available via PyPi or at https://github.com/jgsoltes/hdim-opt, which includes the experimental notebooks for full reproducibility.
-6. Conclusion
+The open-source package containing QUASAR, hdim_opt, is available at https://github.com/jgsoltes/hdim-opt, which includes the experimental notebooks for full reproducibility.
+6 Conclusion
 This study introduced Quasi-Adaptive Search with Asymptotic Reinitialization (QUASAR), a quantum-inspired evolutionary algorithm designed to efficiently explore and exploit complex high-dimensional spaces.
@@
 The empirical evidence validates QUASAR as a powerful evolutionary algorithm, ready for the next generation of high-dimensional optimization problems.
-7. Future Works
-7.1. Enhancements
-QUASAR’s quasi-adaptive mechanisms are highly effective; future improvements will explore fully self-adaptive strategies, similar to those found in other modern DE variants.
-7.1.1. Mutations:
+7 Future Works
+QUASAR’s quasi-adaptive mechanisms may be improved via strategies found in other modern DE variants.
+7.1 Enhancements
+7.1.1 Mutations:
 The mutation factors are sampled from fixed Gaussian distributions; dynamic scaling based on convergence may significantly help to maintain a healthy balance of exploration and exploitation.
 The strategies may benefit from different permutations of equation terms, or non-probablistic selction methods.
-7.1.2. Crossover:
+7.1.2 Crossover:
 Crossover rate is currently defined by fitness rank, and may benefit from a probability-based approach similar to the other QUASAR components, via soft-max or similar methods.
 Many techniques from existing evolutionary frameworks may prove beneficial to the model.
-7.1.3. Reinitialization:
+7.1.3 Reinitialization:
 The current reinitialization covariance matrix ($\mathbf{\Sigma}$) treats all top solutions equally; future work will investigate weighted solutions similar to those found in CMA-ES.
@@
 Additionally, the reinitialization is currently bypassed for problems where $N<D$. Testing other redistribution methods, such as uniform quasi-Monte Carlo sequences, may be beneficial for these cases.
-7.1.4. Verification:
+7.1.4 Verification:
 While QUASAR’s highly stochastic framework makes it a practical choice for non-differentiable and non-parametric problems, the algorithms’s performance would be further validated by rigorously evaluating its performance on constrained or multi-objective problems.
-Acknowledgements.
+Acknowledgements
 The author would like to thank Dr.
@@
 Monet remains a source of inspiration, as always.
-This research was conducted independently and was not part of a formal project.
+This research was conducted independently and was not part of a funded project.